### v1

# A Metamorphic Testing Perspective on Knowledge Distillation for Language Models of Code: Does the Student Deeply Mimic the Teacher?

Abstract
Transformer-based language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, but their practical deployment remains limited due to high computational costs, slow inference speeds, and significant environmental impact. To address these challenges, recent research has increasingly explored knowledge distillation as a method for compressing a large language model of code (the teacher) into a smaller model (the student) while maintaining performance. However, the degree to which a student model deeply mimics the predictive behavior and internal representations of its teacher remains largely unexplored, as current accuracy-based evaluation provides only a surface-level view of model quality and often fails to capture more profound discrepancies in behavioral fidelity between the teacher and student models. To address this gap, we empirically show that the student model often fails to deeply mimic the teacher model, resulting in up to 285% greater performance drop under adversarial attacks, which is not captured by traditional accuracy-based evaluation. In addition, to capture discrepancies in behavioral fidelity, we propose MetaCompress, a metamorphic testing framework that systematically evaluates behavioral fidelity by comparing the outputs of teacher and student models under a set of behavior-preserving metamorphic relations. We evaluate MetaCompress on two widely studied tasks, clone detection and vulnerability prediction, using compressed versions of popular language models of code, CodeBERT and GraphCodeBERT, obtained via three different knowledge distillation techniques: Compressor, AVATAR, and MORPH. The results show that MetaCompress identifies up to 62% behavioral discrepancies in student models, underscoring the need for behavioral fidelity evaluation within the knowledge distillation pipeline and establishing MetaCompress as a practical framework for testing compressed language models of code derived through knowledge distillation.
keywords:
Large Language Models of Code, Knowledge Distillation, Adversarial Attack, Metamorphic Testing, Behavioral Fidelity
[…]
[1]organization=Department of Computer Science, University of Saskatchewan,addressline=110 Science Pl, city=Saskatoon, postcode=S7N 5C9, state=Saskatchewan, country=Canada
1 Introduction
Transformer-based vaswani2017attention language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, including clone detection arshad2022codebert, code summarization ahmed2024automatic, vulnerability prediction ding2024vulnerability, and code search chen2024code. However, despite their impressive capabilities, the adoption of these models remains limited, particularly in consumer-grade devices (e.g., laptops) and latency-sensitive applications, due to their substantial computational demands, slow inference speed, and considerable energy consumption and carbon footprint shi2022compressing, schwartz2020green. For instance, CodeBERT, a widely used code language model with 125 million parameters and a size of 476 MB, can incur up to 1.5 seconds of latency, consume 0.32 kWh of energy, and emit 0.14 kg of CO2 on consumer-grade laptops. [^1] When integrated into IDEs, CodeBERT may be invoked thousands of times daily by a developer, reflecting typical usage patterns hellendoorn2019code. To address these practical challenges, the software engineering research community has recently begun exploring knowledge distillation methods hinton2015distilling to develop extremely compact models that reduce model size, lower inference latency, and mitigate environmental impact while maintaining task performance shi2022compressing, shi2024greening, shi2024efficient, wei2023towards, saad2024alpine, d2024compression, chen2025smaller.
Knowledge distillation (KD) transfers knowledge from a large teacher model to a smaller student [^2] In this study, the term student model refers to the compressed model obtained through knowledge distillation, while the term teacher model denotes the original uncompressed model. model, maintaining comparable accuracy while requiring significantly fewer computational resources. In this context, Shi et al. shi2024greening introduced Compressor, a genetic algorithm-based approach that compresses code-specialized language models (e.g., CodeBERT feng2020codebert) from 476 MB and 125 million parameters to only 3 MB with minimal loss in accuracy. Although KD effectively reduces computational costs and environmental impact, student model performance is typically evaluated using traditional accuracy-based metrics that assess whether the student model’s task performance on ground-truth labels remains comparable to that of the teacher model. However, this form of evaluation offers only a surface-level view and may overlook critical discrepancies in behavioral fidelitythe extent to which a student model deeply mimics the predictive behavior and internal representations of its teacher counterpart.
We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks yang2022natural, zhang2020generating, zeng2022extensive that rename identifiers based on the contextual semantics of each code snippet. The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model. This finding indicates that the student model does not deeply mimic the teacher’s predictive behavior and internal representations, highlighting the limitations of the traditional accuracy-based evaluation, which often overlooks more profound behavioral discrepancies between teacher and student models. It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
To fill this gap, we introduce MetaCompress, a testing framework inspired by metamorphic testing principles in software engineering chen2018metamorphic, to evaluate the behavioral fidelity of the student model in knowledge distillation. In contrast to traditional accuracy-based evaluation, which compares model performance against ground-truth labels, MetaCompress performs a head-to-head comparison of student and teacher model outputs across a suite of behavior-preserving metamorphic relations, enabling a more nuanced assessment of behavioral fidelity. We design four metamorphic relations, grounded in strong empirical evidence from manual investigations, that capture discrepancies in the behavioral fidelity of student models, ranging from prediction-level disagreement to divergences in the probability distribution.
We empirically evaluate MetaCompress on two widely studied software engineering tasks, clone detection and vulnerability prediction, using two popular language models of code, CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert. To distill their smaller student models, we apply three state-of-the-art knowledge distillation techniques designed for language models of code: Compressor shi2022compressing, AVATAR shi2024greening, and MORPH panichella2025metamorphic. Our extensive experiments show that MetaCompress can reveal behavioral fidelity discrepancies of up to 62% across four metamorphic relations that traditional accuracy-based metrics fail to detect.
In summary, the main contributions of this paper are as follows:
  • Insight: We are the first to demonstrate that traditional accuracy-based metrics fall short in capturing behavioral fidelity between student and teacher models by measuring adversarial robustness. To address this, we reframe this evaluation challenge as a classical software testing problem and address it through the lens of metamorphic testing. Ensuring behavioral fidelity is critical because it guarantees that the student model not only provides correct outputs but also behaves consistently across scenarios, just like the teacher model, which is essential for complex tasks like coding.
  • Technique: We propose MetaCompress, a novel output-based metamorphic testing framework grounded in four behavior-preserving metamorphic relations to systematically evaluate the behavioral fidelity of student models for code
  • Evaluation: We conduct a comprehensive evaluation of MetaCompress, demonstrating its effectiveness in uncovering discrepancies in the behavioral fidelity of student models for code that are often overlooked by traditional accuracy-based metrics.
  • Open science: Our replication package has been open-sourced for transparency and reproducibility [^3] https://doi.org/10.5281/zenodo.16127320.
2 Preliminaries
2.1 Language Models of Code
Transformer-based vaswani2017attention language models of code have demonstrated remarkable performance across a wide range of software analytics tasks, such as vulnerability prediction, clone detection, code summarization, code completion, code comment generation, and code search feng2020codebert, guo2020graphcodebert, lu2021codexglue, ahmad2021unified, wang2021codet5, devlin2018bert, liu2019roberta, sanh2019distilbert, husain2019codesearchnet, hou2024large. These models learn general language representations from large amounts of unlabeled data through a two-step process: (i) pre-training and (ii) fine-tuning. During pre-training, the model employs a self-supervised objective to acquire a general understanding of source code structures and patterns. Once the training is complete, the model can be fine-tuned by updating a few parameters for specific tasks with smaller datasets, making it more efficient in terms of time and computational resources compared to training a model from scratch devlin2018bert.
Language models for code can be broadly classified into three categories: (1) encoder-only, (2) decoder-only, and (3) encoder-decoder models zeng2022extensive. Encoder-only models (e.g., CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert) process the entire input simultaneously and are primarily used for discriminative tasks, such as classification (e.g., vulnerability prediction and clone detection). Decoder-only models (e.g., CodeGPT lu2021codexglue, StarCoder lozhkov2024starcoder, and Codex chen2021evaluating) generate output token by token and are typically employed in generative tasks, such as code generation, code completion, and summarization. Encoder-decoder models (e.g., PLBART ahmad2021unified and CodeT5 wang2021codet5) encode the input into a latent representation and subsequently decode it into an output sequence, making them suitable for translation-style tasks such as code summarization and programming language translation. Among these categories, encoder only models, particularly CodeBERT and GraphCodeBERT, have been widely and effectively applied to various downstream software analytics tasks and model compression studies shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer, as evidenced by prior empirical evaluations lu2021codexglue. Consequently, these models serve as the primary focus of our study.
2.2 Model Compression
A significant barrier to the real-world deployment of language models of code is their substantial computational overhead, high memory demands, and environmental impact (e.g., carbon emissions) schwartz2020green, xu2023survey. To mitigate these challenges and advance sustainability, a widely adopted strategy known as model compression offers a solution that reduces the resource consumption of such models without significantly compromising performance. Model compression refers to the process of reducing a large, computationally intensive model into a smaller, more efficient version suitable for deployment on resource-constrained devices (e.g., consumer-grade laptops). In addition to deployment benefits, model compression improves inference speed and enhances overall resource efficiency. Algorithmically, the fundamental model compression techniques include: pruning sanh2020movement, quantization zafrir2019q8bert, knowledge distillation hinton2015distilling, neural architecture search elsken2019neural, weight sharing & hashing chen2015compressing, and low-rank factorization sainath2013low-each method uniquely contributes to model optimization. Among these, pruning, quantization, and knowledge distillation (KD) have emerged as the most prominent and widely adopted approaches for compressing language models of code in recent software engineering research. shi2022compressing, shi2024greening, wei2023towards, d2024compression, chen2025smaller, liu2026pioneer, afrin2025quantization.
While pruning and quantization primarily remove redundant parameters or reduce numerical precision to achieve compactness, their ability to produce highly compact models is inherently limited, as aggressive compression often results in substantial degradation of representational capacity. In contrast, KD offers a distinct advantage by enabling the creation of a smaller student model that can closely approximate, and in some cases replicate, the predictive behavior of its larger teacher counterpart. Through this process, the student learns to mimic the teacher’s functional behavior using soft targets such as probability distributions, internal embeddings, or logits generated by the teacher. This allows the student model to preserve the teacher’s predictive behavior and generalization while maintaining a significantly reduced computational footprint. In the context of language models of code, KD has emerged as a promising compression technique for producing a deployable model that retains functional equivalence with its teacher, although the extent to which a student model truly mimics its teacher’s behavior remains largely unexplored in software engineering.
2.3 Metamorphic Testing
Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem chen2018metamorphic, zhang2014search. Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs). An MR specifies a necessary property that captures the expected relationship between multiple inputs and their corresponding outputs. These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified. This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed yang2025hallucination, xie2011testing, chen2020metamorphic, xiao2022metamorphic, segura2018metamorphic.
In this study, we adapt MT to evaluate how deeply a student model produced via KD mimics its teacher model’s behavior. Unlike conventional MT, we do not apply transformations to the input samples. Instead, we define MRs between the teacher and student models’ outputs, ensuring that when both receive the same input, their outputs exhibit a consistent relationship. These teacher–student MRs capture multiple aspects of behavioral fidelity, including agreement in predicted labels, similarity in probability distributions, and calibration consistency. By systematically evaluating these MRs, we gain a deeper understanding of behavioral fidelity between the teacher and student models beyond accuracy-based evaluation.
2.4 Adversarial Attack
An adversarial attack refers to the deliberate process of crafting adversarial examples by introducing subtle perturbations to the original inputs. Although these perturbed instances appear almost identical to their original counterparts from a human perspective, they can cause the model to yield erroneous or inconsistent outputs. The concept of adversarial attack was first introduced by Zegedy et al. szegedy2013intriguing, who demonstrated that state-of-the-art image classifiers can be misled by carefully designed pixel-level modifications that remain imperceptible to human observers. Formally, given a classifier $f:X\rightarrow Y$ that correctly predicts the label $y_{truth}\in Y$ for an input $x\in X$, an adversarial attack generates a perturbed version $x^{\prime}$ such that $f(x^{\prime})\neq f(x)=y_{truth}$.
In continuous domains such as images, producing visually indistinguishable adversarial examples is relatively straightforward. However, generating adversarial examples for discrete domains like source code is considerably more complex due to the rigid syntactic and grammatical structure of programming languages zhang2022towards, zhang2020generating. As a result, adversarial examples designed to attack language models of code must meet three fundamental requirements. First, the perturbation must effectively mislead the model and reduce its predictive performance. Second, the perturbed code must remain syntactically correct, adhering to the formal rules of the programming language; for example, in Python, identifiers cannot begin with digits and may only contain letters, digits, or underscores. Third, the perturbed code $x^{\prime}$ must preserve the semantics of the original code $x$, maintaining identical functionality and behavior when executed under the same conditions. When an input instance is modified in accordance with these constraints to generate a perturbed version, the resulting instance constitutes an adversarial example; if it causes the model to produce an incorrect or inconsistent output, it is regarded as an adversarial attack.
2.5 Adversarial Robustness
Adversarial robustness describes a model’s capacity to maintain its predictive performance and stability when exposed to adversarial perturbations. A model is considered adversarially robust if it continues to generate consistent, accurate outputs despite deliberate input modifications intended to mislead it. In the context of machine learning, this concept reflects a model’s resilience to adversarial attacks that exploit its sensitivity to small perturbations in the input space szegedy2013intriguing, zhang2022towards.
2.6 Kullback–Leibler Divergence
The Kullback-Leibler (KL) divergence kullback1951information is a measure of how one probability distribution $Q$ diverges from a reference probability distribution $P$. For discrete distributions defined over the same support, it is mathematically expressed as:
$D_{\mathrm{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
where $P(i)$ and $Q(i)$ represent the probabilities of the $i$-th event under distributions $P$ and $Q$, respectively. KL divergence is always non-negative and equals zero if and only if $P=Q$. It is commonly used in machine learning to quantify the difference between predicted and target distributions, particularly in tasks such as knowledge distillation. An important characteristic of KL divergence is that it is not symmetric, meaning that in general, $D_{\mathrm{KL}}(P\parallel Q)\neq D_{\mathrm{KL}}(Q\parallel P)$.
3 Empirical Study
This section presents the design of our empirical study, which aims to evaluate whether the student model deeply mimics the teacher’s predictive behavior and internal representations during knowledge distillation by assessing its adversarial robustness. We hypothesize that if the student model deeply mimics the teacher during knowledge distillation, it should exhibit adversarial performance comparable to the teacher’s. Any substantial deviation in robustness would indicate that the student model fails to deeply mimic the teacher’s predictive behavior and internal representations. We first describe the selected knowledge distillation approaches for code-language models, followed by details on datasets, models, adversarial attack approaches, and the experimental setup. Finally, we outline the evaluation metrics used in the analysis and present the corresponding results in the following section.
3.1 Study Design
3.1.1 Knowledge Distillation Approaches
In recent years, several knowledge distillation (KD) approaches have been proposed for language models of code, including Compressor shi2022compressing, AVATAR shi2024greening, SODA chen2025smaller, MORPH panichella2025metamorphic, and PIONEER liu2026pioneer. In this study, we select Compressor, AVATAR, and MORPH as our KD approaches to generate student models, as the replication package for SODA is not fully available [^4] https://shorturl.at/Se0f7. Furthermore, both MORPH and PIONEER incorporate robustness as an additional objective of knowledge distillation; however, because MORPH optimizes a broader set of objectives than PIONEER, we exclude PIONEER from our empirical investigation.
Compressor, AVATAR, and MORPH generate deployable models from language models of code using neural architecture search guided by multi-objective optimization. Compressor explores various architecture-related hyperparameters, including the number of network layers, dimensionality of network layers, number of attention heads, dimensionality of the feed-forward layers, and vocabulary size. Whereas, both AVATAR and MORPH use the same configuration space for distillation. The configuration parameters include the choice of tokenizer (BPE, WordPiece, Unigram, or Word), a vocabulary size ranging from 1,000 to 46,000, and number of hidden layers ranging from 1 to 12. Hidden layer dimension varies from 16 to 256, with activation functions selected from gelu, relu, silu, and gelu_new. Hidden and attention dropout probabilities are each set to a value between 0.2 and 0.5. Intermediate layer size ranges from 32 to 3,072, and number of attention heads ranges from 1 to 12. Maximum sequence length is configured between 256 and 512, while position embedding type can be absolute, relative_key, or relative_key_query.
Compressor uses the genetic algorithm to identify a simplified architecture that maximizes giga floating-point operations (GFLOPs) while minimizing model size within a fixed memory budget. AVATAR compresses language models of code into deployable versions by jointly optimizing model size, inference latency, energy consumption, and accuracy. It formulates the problem as a multi-objective configuration tuning task, using an SMT solver to define the search space and a tailored optimization algorithm to identify a Pareto-optimal set of configurations for the student model. Based on the AVATAR, MORPH considers metamorphic robustness a crucial property for multi-objective optimization in architecture selection for the student model. In addition, it distills knowledge from the teacher model not only on the soft targets but also on the metamorphic code’s outputs, assuming the teacher model always produces correct outputs for the metamorphic code. Subsequently, knowledge distillation is performed by minimizing the loss between the soft targets produced by the teacher model and those generated by the student model. For example, Compressor employs the KL divergence loss function, defined as follows:
$\mathcal{L}=-\frac{1}{n}\sum_{i}^{n}softmax(\frac{p_{i}}{T})\text{log}\Bigl(softmax(\frac{q_{i}}{T})\Bigl)T^{2}$ (2)
In “(2)”, $p_{i}$ and $q_{i}$ represent the teacher and student model logits, respectively, with $T$ denoting the temperature parameter controlling the softmax distribution hinton2015distilling. While the teacher logits remain fixed, the student logits are updated throughout the training process. It is essential to note that we utilize off-the-shelf compressed versions of CodeBERT and GraphCodeBERT, produced by MORPH and AVATAR, for clone detection and vulnerability prediction tasks in this empirical evaluation.
3.1.2 Datasets and Subjects
To minimize experimental bias in evaluating whether the student model deeply mimics the teacher’s predictive behavior and internal representations during knowledge distillation, we adopt the same models and datasets used in the knowledge distillation–based approaches described in Section 3.1.1. Specifically, we select CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert models and two well-studied downstream tasks: clone detection and vulnerability prediction, which serve as standard benchmarks in knowledge distillation research for software engineering shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer.
Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent. To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset svajlenko2014towards provided by Wang et al. wang2020detecting. The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing. Vulnerability prediction: (VP) This task aims to determine whether a given code snippet contains a vulnerability. We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al. zhou2019devign. The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu. Each function is manually annotated to indicate whether it contains vulnerabilities. We use the CodeXGLUE benchmark splits lu2021codexglue, including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa liu2019roberta. It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768. CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input. The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset husain2019codesearchnet. GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training. Additionally, it introduces a pre-training task that predicts whether a data flow edge exists between two nodes.
3.1.3 Adversarial Attack Approaches
In this study, adversarial attacks are employed to evaluate how deeply the student model mimics the teacher’s predictive behavior and internal representations during knowledge distillation. Adversarial attacks are typically categorized as black-box or white-box, depending on the attacker’s level of access to the model. Following the recommendation of Du et al. du2023extensive, we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code yang2022natural, zhang2020generating. The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions. To this end, we employ three established black-box techniques: ALERT yang2022natural, Metropolis Hastings Modifier (MHM) zhang2020generating, and WIR-Random zeng2022extensive, each of which performs identifier substitution to assess model robustness without altering code functionality. ALERT yang2022natural considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier. MHM zhang2020generating utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration. WIR-Random zeng2022extensive ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates. It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications yang2022natural, zhang2020generating.
3.1.4 Experimental Settings
We fine-tune CodeBERT using the hyperparameter settings and training pipeline provided by the CodeXGLUE benchmark lu2021codexglue, while for GraphCodeBERT, we adopt the settings specified in its original paper guo2020graphcodebert. To ensure consistency across our experiments, we adopt the data-splitting strategy used in knowledge distillation–based approaches such as MORPH and AVATAR, which divide the training data into two equal halves: one for fine-tuning and the other for knowledge distillation. We apply the same fine-tuning split to all models. Despite being trained on only half of the original dataset, all models achieve performance comparable to the state-of-the-art results reported in the CodeXGLUE benchmark lu2021codexglue, the GraphCodeBERT paper guo2020graphcodebert, and knowledge distillation–based compression studies such as Compressor, AVATAR, and MORPH.
3.1.5 Evaluation Metrics
Following the evaluation protocols of CodeXGLUE lu2021codexglue, Yang et al. yang2022natural, and Shi et al. shi2022compressing, shi2024greening, we use accuracy as the primary metric to evaluate model performance on downstream tasks.
To evaluate the performance of both teacher and student models under adversarial conditions, we measure the effectiveness of the applied attack methods. A higher attack effectiveness indicates greater model vulnerability to adversarial perturbations. We employ the Attack Success Rate (ASR) metric yang2022natural to measure the effectiveness of an attack method. ASR measures the proportion of adversarial samples for which the model fails to produce a correct prediction after the attack and is formally defined as follows:
$\text{ASR}=\frac{|\{x|x\in X\wedge M(x^{\prime})\neq M(x)\}|}{|X|}$ (3)
Where $x$ is an original input, $x^{\prime}$ is a modified input. A higher ASR value indicates greater attack effectiveness, reflecting the models’ adversarial vulnerabilities.
The robustness of compressed models to adversarial attacks depends strongly on the quality of the adversarial examples generated. If these examples contain syntactic or semantic errors, model mispredictions may not accurately reflect a lack of robustness but rather the model’s inability to handle unrealistic inputs. Moreover, excessive modifications to the original code during adversarial example generation can lead to out-of-distribution instances, thereby reducing the reliability of robustness assessments. To ensure a fair and meaningful evaluation, we conduct a qualitative analysis to assess the effectiveness of the generated adversarial examples and to verify that the applied perturbations are minimal yet sufficient to mislead the models. Because the inputs for the selected software analytics tasks are source code, we employ four established metrics from Du et al. du2023extensive: Identifier Change Rate (ICR), Token Change Rate (TCR), Average Edit Distance (AED), and Average Code Similarity (ACS), to measure the quality of the adversarial examples.
For a set of $m$ adversarial examples, each $i$-th code snippet contains $k_{i}$ identifiers, among which $n_{i}$ are modified to generate the adversarial example. The ICR is computed as follows:
$\text{ICR}=\frac{\sum_{1}^{m}n_{i}}{\sum_{1}^{m}k_{i}}$ (4)
The TCR represents the proportion of modified tokens relative to the total number of tokens in the original code. The AED quantifies the average number of character-level edits required to transform one token into another, reflecting the textual intensity of the perturbation. The ACS measures cosine similarity between embeddings produced by CodeBERT before and after the adversarial modifications. In general, high-quality adversarial examples are characterized by lower ICR, TCR, and AED values, along with higher ACS scores, indicating that the perturbations are subtle yet effective while preserving code correctness.
3.2 Empirical Results
This section presents the empirical results of our study designed to address the first research question, RQ1: Does the student model deeply mimic the teacher’s predictive behavior and internal representations during knowledge distillation?
We address this research question by evaluating the performance of student models under adversarial attacks relative to their teacher counterparts. In addition, we examine whether traditional accuracy-based evaluation can capture discrepancies in behavioral fidelity arising from the student models’ inability to deeply mimic their teacher models. Since AVATAR outperforms Compressor and MORPH considers metamorphic robustness during knowledge distillation, we compare the adversarial robustness of student models compressed by AVATAR and MORPH with that of the teacher model for this empirical analysis.
To answer this research question, we first compare the performance of the teacher and student models on the original, unperturbed test samples using accuracy as the evaluation metric. Table 1 presents the accuracy comparison between teacher and student models across the clone detection and vulnerability prediction tasks. Overall, the results indicate that the student models maintain performance levels comparable to their teacher counterparts, with accuracy differences remaining within a narrow margin (mostly below 3%). For instance, in clone detection, the student models CodeBERT and GraphCodeBERT compressed by AVATAR and MORPH achieve accuracies close to those of their teacher models, with only minor reductions observed for AVATAR. A similar trend is observed in vulnerability prediction, where student models exhibit minimal degradation and, in some cases, nearly identical performance to the teachers.
Table 1: Accuracy (%) comparison between teacher (uncompressed) and student (compressed) models across clone detection and vulnerability prediction tasks.
Task Model Teacher Compressor AVATAR MORPH
Clone Detection CodeBERT 96.7 95.5 93.1 95.7
GraphCodeBERT 96.5 93.9 93.6 96.3
Vulnerability Prediction CodeBERT 61.7 59.2 60.6 59.4
GraphCodeBERT 60.8 60.3 59.3 59.7
Although the student models achieve task performance comparable to that of their teachers, these results provide only a surface-level view of knowledge transfer. Accuracy-based evaluation captures output correctness but may overlook discrepancies in behavioral fidelity between teacher and student models. Two models may achieve similar accuracy yet differ in representation, generalization, or response to perturbations. Hence, accuracy alone may not reveal whether the student model deeply mimics the teacher’s predictive behavior and internal representations. To address this, we further examine both models under adversarial attack scenarios to gain deeper insight into their behavioral fidelity.
Before analyzing model robustness under adversarial conditions, we first evaluate the quality of the adversarial examples generated by each attack technique. Excessive or unrealistic modifications can distort the syntactic or semantic structure of the original code, leading to misleading robustness assessments. To ensure reliability, we examine the syntactic and semantic integrity of the adversarial examples produced by each attack technique, confirming that the applied perturbations are minimal yet effective. This analysis uses the CodeBERT and GraphCodeBERT models trained on the clone detection task, with complete results for the vulnerability prediction task available in our replication package. Since all attack techniques modify only identifiers, the resulting code retains its original semantics. Figure 1 illustrates the quality assessment of adversarial examples generated by different attack techniques against both teacher and student models.
Figure 1: Quality analysis of adversarial examples across various models and evaluation metrics.
Overall, the results indicate that these attack techniques effectively preserve the syntactic integrity of the generated adversarial examples. Specifically, the examples consistently exhibit high ACS values and low TCR scores across both CodeBERT and GraphCodeBERT models, demonstrating minimal syntactical distortion. The adversarial examples also maintain strong overall quality across the AED and ICR metrics for both the teacher and student models. Although a slight degradation is observed in certain cases—for instance, the MHM attack shows comparatively higher AED values—the overall differences remain limited. We further perform the Friedman test friedman1937use to determine whether the three attack techniques (ALERT, MHM, and WIR-Random) generate adversarial examples of significantly different quality for each model variant of CodeBERT and GraphCodeBERT. A $p$-value below 0.05 indicates that the differences in the quality of adversarial examples produced by the three attack techniques are statistically significant. However, the experimental results show no statistically significant differences among the attack methods for either CodeBERT ($p=0.47$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively) or GraphCodeBERT ($p=0.63$, $p=0.37$, and $p=0.37$ for Teacher, AVATAR, and MORPH, respectively). These results indicate that all three attack techniques generate adversarial examples of comparable quality across the evaluated metrics (ICR, TCR, AED, and ACS).
Although the preceding results show that the adversarial examples essentially preserve syntactic and semantic quality across both teacher and student models, it remains important to examine whether any subtle quality differences exist between these two model categories. To investigate this, we compare the quality of adversarial examples generated by each attack technique for the teacher and student models across all the evaluation metrics. The Wilcoxon signed-rank test woolson2007wilcoxon is used to assess the statistical significance of these differences. The comparison between the Teacher and AVATAR models ($p=0.13$) and between the Teacher and MORPH models ($p=0.85$) indicates that the quality of adversarial examples remains statistically comparable, with no significant differences observed in either case. These findings suggest that the attack techniques generate adversarial examples of similar syntactic and semantic integrity across both teacher and student models, confirming that the subsequent robustness analysis is based on adversarial inputs of consistent quality and that any observed performance degradation can be attributed to the models’ inability to maintain robustness against adversarial perturbations.
Having rigorously evaluated the quality of the generated adversarial examples, we next examine how effectively the teacher and student models withstand adversarial attacks. Figure 2 presents the Attack Success Rate (ASR) values for the teacher and student models (AVATAR and MORPH) across the ALERT, MHM, and WIR-Random attack techniques for both the clone detection and vulnerability prediction tasks. Overall, the results indicate that student models produced through knowledge distillation exhibit substantially lower robustness than their corresponding teacher models, as evidenced by consistently higher ASR values across most attack settings, with only two exceptions. In the clone detection task, the teacher model demonstrates lower robustness than the MORPH student model under the MHM and WIR-Random attacks, and than the AVATAR student model under the MHM attack. Although the robustness differences between the teacher and student models are relatively small for the GraphCodeBERT model and the vulnerability prediction task, they are considerably larger in all other experimental settings. For instance, in the clone detection task with the CodeBERT model under the ALERT attack, the AVATAR student model experiences approximately three times greater performance degradation than the teacher model.
Figure 2: Attack success rate metric values across adversarial attack techniques, models, tasks, and knowledge distillation methods.
Taken together, these findings suggest that student models distilled from their teachers, particularly those produced by AVATAR and, to a lesser extent, by MORPH, are more susceptible to adversarial attacks than their teacher counterparts. Surprisingly, although MORPH generates student models by incorporating metamorphic robustness during knowledge distillation, the resulting student models exhibit lower robustness than those produced by AVATAR across all the attack techniques for the GraphCodeBERT model in the vulnerability prediction task. The results further reveal that student models fail to deeply mimic the teacher’s predictive behavior and internal representations. Although knowledge distillation preserves task accuracy under clean inputs, it weakens the model’s ability to generate stable and consistent predictions when exposed to adversarial perturbations. As a result, traditional accuracy-based evaluation cannot capture these discrepancies in behavioral fidelity, which become evident only under adversarial conditions. These observations highlight the need for more comprehensive evaluation frameworks that assess behavioral fidelity beyond accuracy.
Answer to RQ1 Since the adversarial examples across all teacher and student models exhibit comparable syntactic and semantic quality, the observed robustness gap can be attributed to the student models’ inability to deeply mimic the teacher models’ predictive behavior and internal representations, a limitation that traditional accuracy-based evaluation fails to reveal.
4 Approach: MetaCompress Framework
The empirical findings in Section 3.2 show that traditional accuracy-based evaluation is insufficient to capture behavioral discrepancies between teacher and student models, which become evident under adversarial conditions due to the student models’ inability to deeply mimic their teacher counterparts’ predictive behavior and internal representations. Motivated by this observation, we introduce MetaCompress, a metamorphic testing framework for evaluating the behavioral fidelity of compressed models by framing the task as a classical software testing problem and applying a metamorphic testing strategy. The framework identifies key behavioral invariants that should be preserved after model compression and formalizes a set of output-based metamorphic relations that serve as oracle-free tests to assess behavioral fidelity between teacher and student models. Figure 3 presents the overall design and workflow of the proposed MetaCompress framework.
Figure 3: Workflow of the MetaCompress framework, comparing teacher and student model outputs under behavior-preserving metamorphic relations to assess behavioral fidelity.
4.1 Behavioral Fidelity as a Software Testing Problem
Knowledge distillation transfers knowledge from a large teacher model to a smaller student model, aiming to preserve the student’s performance on the original task. However, our empirical results demonstrate that the student model does not deeply mimic the teacher’s internal behavior, raising an important question: Does the student model behave consistently with its teacher? This concern closely parallels a classic problem in software engineering: verifying that a modified system preserves the intended behavior of its original version. In this context, we treat the student model as a transformed version of the teacher model and frame the evaluation of behavioral fidelity as a software testing problem. From this perspective, the teacher and student models represent two versions of the same system under test (SUT), and our goal is to assess whether behavioral fidelity is maintained after knowledge distillation. Formally, let $M:\mathcal{X}\rightarrow\mathcal{Y}$ denote the teacher model and $M^{\prime}:\mathcal{X}\rightarrow\mathcal{Y^{\prime}}$ its student variant. Given a set of test inputs $\mathcal{T}\subseteq\mathcal{X}$ and a task-specific equivalence relation $\equiv\subseteq\mathcal{Y}\times\mathcal{Y^{\prime}}$, the problem of behavioral fidelity testing is to determine whether:
$\forall x\in\mathcal{T},\quad M(x)\equiv M^{\prime}(x)$ (5)
That is, whether the student model produces outputs that are behaviorally consistent with those of the teacher model across all test cases.
In traditional software engineering, techniques such as regression testing wong1997study, mutation testing jia2010analysis, and metamorphic testing chen2018metamorphic are commonly employed to assess the reliability of modified systems. However, evaluating language models of code presents unique challenges compared to traditional software systems. For example, traditional software typically produces deterministic outputs, while language models of code yield probabilistic outputs, and knowledge distillation may alter output distributions even when top-1 predictions remain unchanged. Additionally, we aim to assess the behavioral fidelity of a student model compared to its teacher counterpart. Although ground-truth labels are available for tasks such as clone detection or vulnerability prediction, our objective is not to directly compare the outputs of the two models against the ground-truth. Instead, we compare their outputs after model execution, treating the teacher model as the behavioral reference. Notably, the teacher model’s behavior is not known until it is executed on test data. We refer to this scenario as an instance of the oracle problem, where the absence of a definitive oracle hinders straightforward correctness verification.
As a result, conventional software testing techniques such as regression testing, which depend on exact output matches, are often too rigid—minor differences in probabilistic outputs or generated sequences do not necessarily indicate functional errors. Mutation testing is also ill-suited, as neither model is inherently faulty. In contrast, metamorphic testing provides a principled approach to overcoming the oracle problem by defining expected relations (as detailed in Section 4.2) between the outputs of the teacher and student models, without requiring explicit ground-truth. These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs. Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models xie2020mettle, sun2024fairness, kwan2024mt, zhang2020machine, wu2025detecting, yang2025hallucination, demonstrating its versatility and effectiveness in oracle-free evaluation scenarios. Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
4.2 Designing Output-Based Metamorphic Relations
MetaCompress is designed to evaluate discrepancies in behavioral fidelity between teacher and student models. Without a head-to-head comparison of teacher and student models on identical inputs, it is difficult to determine whether the student model maintains the teacher model’s behavioral fidelity. For example, a manual inspection of a student GraphCodeBERT model generated using Compressor for the vulnerability prediction task illustrates this limitation. Figure 3(a) shows a case in which both models predict the same class, yet their output probability distributions diverge substantially, with a Kullback–Leibler divergence of 0.6177. In contrast, as shown in Figure 3(b), the teacher and student models yield different class predictions for the same input, revealing behavioral differences that accuracy alone fails to capture. To systematically assess such differences, MetaCompress defines a set of output-based metamorphic relations that encode expected behavioral invariants, such as class agreement and similarity in probability distributions, between teacher and student models. Unlike traditional metamorphic testing, which typically applies transformations to inputs, these relations operate directly on model outputs, enabling a structured and quantitative evaluation of behavioral fidelity. The following subsections describe the design rationale, empirical motivation, and evaluation metrics for each metamorphic relation incorporated into the MetaCompress framework.
Figure 4: Behavioral fidelity discrepancies between the outputs of teacher and student models.
4.2.1 MR1 - Prediction Agreement
Description
Given a set of test inputs $X={x_{1},x_{2},...,x_{n}}$, a teacher model $M_{T}$ and a student model $M_{S}$, for each input $x_{i}\epsilon X$, the prediction of $M_{S}$ should match of $M_{T}$. We define MR1 as follows:
$M_{S}(x_{i})=M_{T}(x_{i})$ (6)
Rationale
This MR captures the most intuitive form of behavioral fidelity: top-1 label consistency. Our empirical observations (as shown in Figure 3(b)) suggest that prediction disagreement may be associated with divergence in model reasoning. Therefore, MR1 serves as a meaningful indicator of behavioral fidelity.
Evaluation metric
To measure this MR, we consider the Label Loyalty metric that quantifies the degree to which this MR holds across all the test instances. It computes the percentage of instances for which the student model’s prediction matches the teacher model’s prediction, analogous to accuracy:
$\text{Loyalty}_{\text{lab}}=\frac{1}{n}\sum_{i=1}^{n}1[M_{S}(x_{i})=M_{T}(x_{i})]$ (7)
Thus, $\text{Loyalty}_{\text{lab}}$ quantitatively measures the success rate of this MR. A higher metric value indicates greater behavioral fidelity preservation in the student model.
4.2.2 MR2 - Probability Distribution Similarity
Description
This MR assesses the similarity of output probability distributions between the two models by evaluating whether, for a given input $x$, the predicted distribution of $M_{S}$ closely aligns with that of $M_{T}$. Given a set of test inputs $X={x_{1},x_{2},...,x_{n}}$, let $P_{T}(y\mid x)$ and $P_{S}(y\mid x)$ represent the probability distributions over the output space $\mathcal{Y}$ produced by $M_{T}$ and $M_{S}$, respectively. We define MR2 as follows:
$\quad D\left(P_{T}(\cdot\mid x),P_{S}(\cdot\mid x)\right)\leq\delta,\quad\forall x\in{X}$ (8)
Where $D(\cdot,\cdot)$ is a divergence metric (e.g., Kullback–Leibler divergence), and $\delta\geq 0$ is a predefined tolerance threshold.
Rationale
Although the teacher and student models may produce the same predicted label, their underlying probability distributions can differ substantially. Our empirical observations (Figure 3(a)) reveal such differences, as measured by the Kullback–Leibler divergence metric (e.g., 0.6177), suggesting that prediction agreement (MR1) alone is insufficient to ensure behavioral fidelity.
Evaluation metric
To quantify the difference between the two probability distributions, we use the Kullback–Leibler divergence, as defined in “(1)”. Given a tolerance $\delta=0.5$ (Inspired by sigmoid-based methods for binary classification lecun2015deep), we then count the test instances for which MR2 holds:
$\text{Loyalty}_{\text{prob}}=\frac{1}{|X|}\sum_{x\in X}\mathbf{1}\!\left[D_{\mathrm{KL}}\!\bigl(P_{T}(\cdot\mid x),P_{S}(\cdot\mid x)\bigr)\leq\delta\right]\qquad$ (9)
A larger $\text{Loyalty}_{\text{prob}}$ indicates that the student model retains the teacher model’s probability-distribution behavior on a greater portion of the test set.
4.2.3 MR3 - High Confidence Preservation
Description
Given a set of test inputs $X={x_{1},x_{2},...,x_{n}}$, $M_{T}$, and $M_{S}$, consider the subset: $\mathcal{X}_{\text{conf}}=\left\{\,x_{i}\in X\;\middle|\;\max M_{T}(x_{i})\geq\tau\right\},$
where $\tau$ is a high-confidence threshold (e.g., $\tau=0.9$). For every $x_{i}\in\mathcal{X}_{\text{conf}}$, the student model should maintain
$\arg\max M_{S}(x_{i})\;=\;\arg\max M_{T}(x_{i})\quad\land\ \quad\max M_{S}(x_{i})\;\geq\;\tau$ (10)
In other words, $M_{S}$ must predict the same top-1 class as $M_{T}$ and do so with confidence at least $\tau$.
Rationale
A Top-1 agreement alone may mask cases where $M_{S}$ is hesitant while $M_{T}$ is decisive. Our manual investigation (Figure 3(a)) revealed that large confidence gaps (e.g., |0.990.51|=0.48|0.99-0.51|=0.48) often coincide with divergent reasoning, even when predictions align. By enforcing both label and confidence alignment on high-certainty inputs, MR3 probes a deeper layer of behavioral fidelity.
Evaluation metric
We introduce the High Confidence Agreement Rate (HCAR) metric to quantify how often this relation holds:
$\text{HCAR}=\frac{1}{|\mathcal{X}_{\text{conf}}|}\sum_{x\in\mathcal{X}_{\text{conf}}}\mathbb{I}\left[\begin{aligned} &\arg\max T(x)=\arg\max S(x)\\ &\land\ \max S(x)\geq\tau\end{aligned}\right]$ (11)
Here $\mathbb{I}[\cdot]$ denotes the indicator function, which returns 1 if the condition is satisfied and 0 otherwise. A higher HCAR indicates that the student model accurately preserves the teacher’s confident and decisive behavior, whereas a lower value suggests reduced fidelity after knowledge distillation.
4.2.4 MR4 - Calibration Alignment
Description
Given a set of test inputs $X={x_{1},x_{2},...,x_{n}}$, $M_{T}$, and $M_{S}$, we partition the confidence interval $[0,1]$ into $B$ equal-width bins. For each bin $i\in\{1,\dots,B\}$, let
$\text{acc}^{\text{orig}}_{i}=\frac{1}{|X_{i}|}\sum_{x\in X_{i}}1\!\bigl[\arg\max M_{T}(x)=y(x)\bigr]$ (12)
$\text{acc}^{\text{comp}}_{i}=\frac{1}{|X_{i}|}\sum_{x\in X_{i}}1\!\bigl[\arg\max M_{S}(x)=y(x)\bigr]$ (13)
where $X_{i}$ is the subset of inputs whose $M_{T}$ confidence $\max M_{T}(x)$ falls into bin $i$ and $y(x)$ is the ground-truth label. The student model is expected to preserve calibration if, for every bin $i$, its empirical accuracy closely matches that of the teacher.
Rationale
Top-1 agreement alone does not indicate whether the student model is well-calibrated. Poor calibration, where predicted probabilities do not reflect true likelihoods, can lead to over- or under-confidence, which adversely affects downstream decision-making. Prior studies shang2024enhancing, ji2024beware on vision and language models demonstrate that compression techniques such as quantization can significantly distort model calibration, even when accuracy remains unaffected. These findings motivate the design of this MR to assess behavioral fidelity through the lens of calibration alignment.
Evaluation metric
We propose the Expected Calibration Alignment (ECA) metric, defined as the mean absolute difference in bin-wise accuracy to assess this behavioral fidelity:
$\mathrm{ECA}\;=\;\frac{1}{B}\sum_{i=1}^{B}\left|\,\text{acc}^{\text{orig}}_{i}-\text{acc}^{\text{comp}}_{i}\right|$ (14)
A lower ECA score ($\text{ECA}\in[0,1]$) indicates that the student model’s calibration behavior closely aligns with that of the teacher model. Conversely, a higher ECA score reflects greater divergence between the two models’ calibration profiles.
Violation rate computation: To compute the violation rate of a metamorphic relation $\text{MR}_{k}$, we use the complement of its hold rate. Let $\text{HoldRate}_{\text{MR}_{k}}\in[0,1]$ denote the proportion of test inputs for which $\text{MR}_{k}$ holds. Then, the violation rate is defined as:
$\text{ViolationRate}_{\text{MR}_{k}}=1-\text{HoldRate}_{\text{MR}_{k}}$ (15)
This formulation enables us to quantify the extent to which a student model deviates from the behavioral invariants defined by each metamorphic relation. Higher violation rates indicate greater behavioral drift from the teacher model.
Expected Behaviour: A student model is considered behavior-preserving if, for every test input, the student model: (i) produces the exact top-1 prediction as the teacher; (ii) retains a probability distribution that closely matches the teacher’s in terms of uncertainty and class ranking; (iii) replicates the teacher’s high-confidence decisions on unambiguous inputs (e.g., $\max M_{T}(x)!\geq!\tau$); and (iv) maintains aligned calibration curves such that the bin-wise accuracy difference $\lvert\text{acc}^{\text{orig}}_{i}-\text{acc}^{\text{comp}}_{i}\rvert$ remains consistently small.
4.3 MetaCompress Evaluation
To evaluate the effectiveness of the proposed MetaCompress framework, we perform an extensive empirical study using the models, datasets, and knowledge distillation techniques described in Section 3.1. This evaluation aims to determine whether MetaCompress can uncover behavioral discrepancies between teacher and student models that traditional accuracy-based metrics fail to capture.
4.3.1 Implementations
MetaCompress is implemented in Python on top of the Compressor, AVATAR, and MORPH. All experiments are conducted on an Ubuntu 22.04 machine equipped with an Intel(R) Xeon(R) W-2265 CPU @ 3.50GHz, 256 GB RAM, and one NVIDIA RTX A6000 48 GB GPU.
4.3.2 Evaluation Results
This section presents the empirical evaluation results of the MetaCompress framework to address the second research question, RQ2: To what extent does MetaCompress uncover behavioral discrepancies that may not be captured by the traditional accuracy metric?
We address this research question by assessing the effectiveness of the MetaCompress framework by measuring the degree of violation across the four designed metamorphic relations. Each violation indicates a behavioral inconsistency between the teacher and student models when exposed to identical inputs, thus quantifying the loss of behavioral fidelity after knowledge distillation. We conduct this analysis across two language models of code (CodeBERT and GraphCodeBERT), two software analytics tasks (clone detection and vulnerability prediction), and three knowledge distillation approaches (Compressor, AVATAR, and MORPH).
MR1 violation: The results for MR1 violation shown in Figure 5 reveal that student models exhibit low violation rates for the clone detection task but considerably higher rates for vulnerability prediction. For clone detection, all three distillation techniques maintain strong label agreement with the teacher, with a maximum violation of 7%, where MORPH achieves the lowest (3%) and AVATAR the highest (7%) inconsistency. In contrast, violation rates rise sharply for vulnerability prediction, ranging from 29% to 36%, indicating notable behavioral divergence between teacher and student models. Among the methods, AVATAR performs slightly better for CodeBERT, while MORPH and Compressor show comparable fidelity across both models. As shown in Table 1, although the student models achieve comparable performance on the vulnerability prediction task (with a maximum accuracy difference of approximately 3%), the MetaCompress framework reveals up to 36% discrepancies in prediction alignment, demonstrating its ability to capture behavioral differences that traditional accuracy-based evaluation overlooks.
Figure 5: MR1 violation rates for different tasks, models, and knowledge distillation techniques.
MR2 violation: The results for MR2 violations shown in Figure 6 reveal that the degree of probabilistic divergence between teacher and student models varies across tasks and distillation techniques. For the clone detection task, the violation rates remain relatively low, ranging from 3% to 13%, suggesting that the student models generally approximate the teacher’s output distributions. In contrast, the vulnerability prediction task shows substantially greater violations, reaching up to 31%, reflecting a pronounced mismatch in the probability distributions of the teacher and student models. These findings suggest that, although the student models can match the teacher’s predictions in terms of label accuracy, their underlying probability distributions differ significantly, underscoring MetaCompress ’s ability to expose fine-grained probabilistic discrepancies beyond traditional accuracy-based evaluation.
Figure 6: MR2 violation rates for different tasks, models, and knowledge distillation techniques.
To gain a finer insight into behavioral fidelity, we visualize the distributions of MR2 violations using box plots, enabling a head-to-head comparison between teacher and student models in terms of output probability deviation across individual test instances. Given the asymmetry of KL divergence, we report both $D_{\mathrm{KL}}(P\parallel Q)$ and $D_{\mathrm{KL}}(Q\parallel P)$, where $P$ and $Q$ are the output probability distributions of the teacher and student models. Figure 7 presents the per-sample distributions of these values for KD approaches: Compressor, AVATAR, and MORPH. We observe that for a large proportion of samples, the KL divergence remains low, indicating close alignment between the teacher and student models. However, a notable number of outliers with high divergence suggest that in some cases, the student model assigns significantly different probabilities to the same output classes. These distributions underscore the utility of MR2 for capturing probabilistic distortions that may not be reflected in the accuracy metric. Together, the distributions of $D_{\mathrm{KL}}(P\parallel Q)$ and $D_{\mathrm{KL}}(Q\parallel P)$ provide a comprehensive understanding of discrepancies in behavioral fidelity at the granularity of individual predictions. These findings demonstrate that MetaCompress effectively exposes subtle differences in probabilistic reasoning between teacher and student models that remain hidden under accuracy-based evaluation.
(a) $D_{\mathrm{KL}}(P\parallel Q)$
(b) $D_{\mathrm{KL}}(Q\parallel P)$
Figure 7: Boxplots of KL divergence illustrating output distribution shifts between teacher and student models across different tasks and knowledge distillation techniques  [^7] CB: CodeBERT, GCB: GraphCodeBERT, VP: Vulnerability Prediction, CD: Clone Detection, AV: AVATAR, CP: Compressor, MP: MORPH, CB-CD-AV: CodeBERT compressed via AVATAR for the Clone Detection task..
MR3 violation: While MR1 and MR2 expose surface-level inconsistencies, such as label disagreement and output distribution shifts, MR3 uncovers deeper deviations in behavioral fidelity. Table 2 illustrates the violations of MR3 observed in student models, as identified by the MetaCompress. Across all tasks and student models, we observe substantial violations of MR3. As shown in Table 2, student GraphCodeBERT models often exhibit significantly lower confidence than their teacher counterparts for the vulnerability prediction task, despite having similar accuracy. For example, the GraphCodeBERT model obtained from AVATAR violates up to 62% of MR3, indicating a significant discrepancy in high-confidence predictions and overall behavioral fidelity between the teacher and student models that accuracy-based evaluation fails to capture. We further examine the sensitivity of MR3 to different threshold values of $\tau$ and find that variations in $\tau$ had a negligible impact on overall violation rates, confirming the metric’s stability.
Table 2: Behavioral fidelity discrepancies in student models revealed through MR3 violations.
[t] Compressor AVATAR MORPH Task Model $\tau=0.8$ $\tau=0.85$ $\tau=0.9$ $\tau=0.8$ $\tau=0.85$ $\tau=0.9$ $\tau=0.8$ $\tau=0.85$ $\tau=0.9$ CB 0.03 0.03 0.03 0.06 0.05 0.04 0.03 0.03 0.03 CD GCB 0.04 0.04 0.04 0.03 0.03 0.03 0.03 0.03 0.03 CB VP GCB 0.44 0.45 0.43 0.6 0.61 0.62 0.5 0.51 0.49
  • Cells highlighted with a dashed line indicate cases where the teacher models do not produce any output with a high confidence score for the given task.
MR4 violation: Table 3 presents the calibration deviation between teacher and student models across three knowledge distillation techniques for varying bin sizes ($B=10,15,20$). The results reveal that the calibration alignment varies notably across tasks and distillation strategies. For the clone detection task, all student models maintain near-perfect calibration consistency with their teachers, with deviation values below 0.04, indicating that the students effectively preserve the teachers’ confidence distributions. In contrast, the vulnerability prediction task exhibits substantially higher deviations marked by dark gray cells, up to 0.14 for CodeBERT and 0.11 for GraphCodeBERT, demonstrating that the student models’ probabilistic calibration diverges considerably from that of the teachers. These results highlight that although student models may achieve comparable accuracy, their confidence calibration can differ sharply, underscoring MetaCompress ’s ability to capture nuanced behavioral discrepancies beyond what traditional evaluation metrics can reveal. We further examine the sensitivity of MR4 to different bin sizes and observe that variations in $B$ had a negligible effect on overall violation rates, confirming the metric’s stability.
Table 3: Behavioral fidelity discrepancies in student models revealed through MR4 violations.
Compressor AVATAR MORPH
Task Model B = 10 B = 15 B = 20 B = 10 B = 15 B = 20 B = 10 B = 15 B = 20
CB 0.017 0.017 0.017 0.035 0.035 0.035 0.002 0.002 0.002
CD GCB 0.026 0.026 0.026 0.063 0.063 0.063 0.003 0.003 0.003
CB 0.136 0.142 0.143 0.016 0.016 0.017 0.133 0.138 0.139
VP GCB 0.092 0.092 0.092 0.112 0.112 0.112 0.094 0.096 0.098
Answer to RQ2 MetaCompress provides a comprehensive evaluation of behavioral fidelity in language models for code, revealing substantial discrepancies that remain hidden under accuracy-based evaluation. Even when the accuracy difference between teacher and student models is within 3%, the framework identifies up to 62% behavioral fidelity violations across the four metamorphic relations. By analyzing model behavior from label agreement (MR1) to probabilistic divergence (MR2), confidence consistency (MR3), and calibration alignment (MR4), MetaCompress captures discrepancies across multiple behavioral dimensions. These findings confirm that MetaCompress effectively uncovers behavioral differences between teacher and student models that conventional accuracy metrics fail to reveal.
5 Discussion
Ba et al. ba2014deep stated that ”If the student model learns to mimic the teacher model perfectly, it makes exactly the same predictions and mistakes as the teacher model.” However, achieving such mimicry is extremely challenging when there exists a substantial gap in network capacity wang2021knowledge. For instance, the knowledge distillation technique Compressor reduces the size of the CodeBERT model from 476 MB to only 3 MB. Consequently, although the student models produced by Compressor, AVATAR, and MORPH achieve accuracy comparable to the teacher, their performance deteriorates considerably under adversarial attacks. Even though MORPH attempts to distill knowledge more robustly by incorporating metamorphic attacks through metamorphic code transformations, the resulting student models still exhibit notable adversarial vulnerabilities. These observations indicate that the knowledge distillation strategies employed by Compressor, AVATAR, and MORPH do not enable the student models to deeply mimic the teacher’s behavior, leaving a substantial knowledge gap in behavioral fidelity between the teacher and the student models.
One might argue that a student model performs comparably to its teacher after knowledge distillation. Then, why should we be concerned about the knowledge gap in behavioral fidelity? We address this question through two key arguments.
  1. 1.
    If the knowledge gap can be reduced by enabling the student model to more deeply mimic the teacher model, then the student should exhibit comparable robustness to its teacher under adversarial conditions. This would make the student model suitable for deployment in real-world software systems where safety and reliability are critical.
  2. 2.
    Wang et al. wang2025empirical empirically demonstrated that the same student model distilled from a teacher performs significantly better than when trained using the standard fine-tuning approach on the same task. Since the teacher model typically possesses much greater network capacity, it can learn more complex decision boundaries and have greater robustness against adversarial attacks. Minimizing the knowledge gap, therefore, enables the distilled student model to better approximate these boundaries during distillation, potentially enhancing its resilience to adversarial attacks.
The empirical findings from our MetaCompress framework substantiate these arguments. Although the accuracy difference between teacher and student models remains within 3%, MetaCompress reveals violations of up to 62% across the four metamorphic relations, indicating substantial discrepancies in behavioral fidelity. For instance, the high MR1 and MR2 violation rates demonstrate that the student models often diverge from the teacher in both label agreement and probability distribution, even when their predictions appear similar at the aggregate level. Likewise, MR3 and MR4 violations expose inconsistencies in confidence and calibration alignment, underscoring that accuracy alone cannot fully capture behavioral divergence. These results confirm that reducing the knowledge gap in behavioral fidelity is essential for developing distilled models that preserve the robustness and decision consistency of their teacher counterparts. These insights open promising avenues for advancing knowledge distillation methods that achieve greater alignment between student and teacher models in terms of predictive behavior, internal representations, generalizability, and robustness.
6 Threats to Validity
This section provides a brief overview of the construct and the internal and external threats associated with our study.
6.1 Internal Validity
A potential threat to internal validity may arise if differences in performance between the student and teacher models are influenced by inconsistent training procedures, fine-tuning configurations, or data handling, rather than the knowledge distillation approaches themselves. To mitigate this threat, we applied uniform training protocols and hyperparameter settings across all models, both student and teacher, following the configurations reported in prior knowledge distillation studies shi2022compressing, shi2024greening, panichella2025metamorphic. Additionally, we used identical data splits for fine-tuning and evaluation to ensure consistency and fairness in comparison. Maintaining such a consistent experimental setup enhances the internal validity of our findings by ensuring that the observed effects are due to the student model’s inability to deeply mimic the teacher model’s behavior, rather than differences in hyperparameters or data partitions. Additionally, the experimental results may be influenced by potential errors in implementing the training pipelines for both student and teacher models. To mitigate this threat, we adopted the standardized training procedures provided by the CodeXGLUE benchmark for teacher model fine-tuning and evaluation. For knowledge distillation, we utilized publicly available, off-the-shelf student models obtained from Compressor [^8] https://github.com/soarsmu/Compressor, AVATAR [^9] https://github.com/soarsmu/Avatar, and MORPH [^10] https://github.com/apanichella/morph-distillation/tree/main.
6.2 Construct Validity
A potential threat to construct validity lies in the assumption that the four metamorphic relations accurately capture behavioral fidelity. To mitigate this, the relations were not defined arbitrarily; instead, they were grounded in empirical observations of how student models diverge from their teacher counterparts on identical inputs (Figure 4). These observed patterns informed the design of each relation, ensuring that they capture behaviorally meaningful differences between student and teacher models. A potential threat to construct validity arises from framing behavioral fidelity evaluation as an oracle problem, which assumes the absence of ground-truth labels despite the availability of test inputs. To mitigate this, we argue that the outputs of the student and the teacher models are not known until execution, much like the behavior of a software system is observed only at runtime. This reinforces the validity of framing behavioral fidelity evaluation as an oracle problem. Another potential threat to construct validity arises from the inherent randomness in model behavior, which can lead to misleading conclusions about behavioral differences. To address this issue, we fixed the random seeds and conducted head-to-head comparisons between each student model and its teacher counterpart. The comparisons used identical inputs across all four metamorphic relations to ensure that any observed differences were due to the student model’s inability to deeply mimic the teacher model’s behavior rather than random variation.
6.3 External Validity
A potential external threat is the generalizability of MetaCompress beyond program understanding to tasks such as code summarization. This study focuses exclusively on knowledge distillation-based approaches, including Compressor, AVATAR, and MORPH, which are designed for code-language models and operate on the teacher model’s logits. Since these methods compute the loss between teacher and student logits for knowledge distillation, they are not directly applicable to generation tasks, where it is uncommon to define a loss function over generated sequences that typically differ in length between the teacher and the student models. This limitation represents an inherent constraint of the knowledge distillation approaches considered in this study. Nevertheless, because MetaCompress evaluates behavioral fidelity based on model outputs rather than internal architectures, it remains applicable to generation tasks once knowledge distillation techniques are developed for those scenarios. The choice of models, tasks, and knowledge distillation approaches may limit the generalizability of our results. To mitigate this, we employed widely adopted models (CodeBERT, GraphCodeBERT), standard tasks (clone detection, vulnerability prediction), and representative knowledge distillation techniques (e.g., Compressor, AVATAR, and MORPH). Future work may extend our analysis to additional models and tasks. To facilitate such efforts, we have thoroughly documented our framework design and evaluation setup and publicly released all code and scripts used in our experiments.
7 Related Work
A substantial body of research on model compression has emerged in the fields of computer vision and natural language processing (NLP), aiming to develop large language models that are energy efficient, environmentally sustainable, and optimized for memory usage xu2023survey, zhu2024survey, xu2025resource, xu2021survey, gordon2020compressing, xu2021beyond, ye2019adversarial, du2021robustness, gourtani2024improving, zhu2022safety, sanh2019distilbert, sun2019patient, jiao2019tinybert, buciluǎ2006model, jiang2023lion, zhang2018structadmm, tang2019distilling, xu2020bert. To this end, several techniques, including pruning sanh2020movement, quantization zafrir2019q8bert, and knowledge distillation (KD) hinton2015distilling, have been widely explored. More recently, these compression techniques have been extended to software analytics to improve the efficiency of language models of code. Wei et al. wei2023towards evaluated quantized models on code generation tasks, examining accuracy, resource consumption, and environmental impact, and found that quantization can substantially improve model efficiency with minimal degradation in accuracy or robustness when applied under suitable conditions. Saad et al. saad2024alpine introduced ALPINE, a pruning method that is independent of programming language and effectively reduces the computational cost of code-language models. Aloisio et al. d2024compression conducted an empirical study on how different compression strategies, such as pruning, quantization, and knowledge distillation, affect the performance and resource efficiency of the CodeBERT model across three software analytics tasks: code summarization, vulnerability detection, and code search. Afrin et al. afrin2025quantization empirically examined the effects of quantizing large code models beyond functional correctness, focusing on qualitative aspects and static properties of the generated code. While pruning and quantization improve efficiency, their compression capacity is limited, as excessive compression often harms model performance. Hence, this study focuses on knowledge distillation, which enables extremely compact models without compromising significant predictive performance.
Shi et al. shi2022compressing proposed Compressor, a KD-based method evaluated on CodeBERT and GraphCodeBERT for two code understanding tasks: vulnerability prediction and clone detection. Experimental results show that Compressor significantly reduces inference time and model size, down to 3 MB, while preserving model effectiveness. In a follow-up study, Shi et al. shi2024greening introduced AVATAR, another KD-based approach designed to mitigate energy consumption and carbon emissions, while also maintaining model effectiveness and reducing inference latency. Building on AVATAR, MORPH panichella2025metamorphic was proposed as a KD method that combines metamorphic testing with many-objective optimization to enhance model robustness on metamorphic code, as well as accuracy, size, and efficiency. Experiments on CodeBERT and GraphCodeBERT showed that MORPH achieved up to 73% greater robustness and 38% greater efficiency than AVATAR, without sacrificing accuracy or increasing model size. However, in MORPH, the metamorphic attack using a metamorphic sample technically differs from the adversarial attack using an adversarial example. Metamorphic attacks are performed only once on the metamorphic sample, whereas adversarial attacks are performed by repeatedly querying the model, modifying the original sample to assess the model’s actual robustness. Our empirical evaluation also reveals that the student models obtained from MORPH experience greater performance degradation under adversarial attacks than under metamorphic attacks, compared to the student models obtained from AVATAR.
Recently, Liu et al. liu2026pioneer proposed a KD-based approach, PIONEER, that enhances the robustness of compressed code models while preserving accuracy. Using a solver-constrained genetic algorithm and robustness evaluation with ALERT-generated examples, it achieved an 87.54% robustness gain with only a 1.67% drop in accuracy across CodeBERT, GraphCodeBERT, and CodeT5. Recently, Wang et al. wang2025empirical empirically evaluated the effectiveness of logit-based and feature-based KD approaches for code understanding tasks. Their study showed that KD consistently outperforms fine-tuning on code understanding tasks, with feature-based methods and code-specific teachers yielding the best results.
Despite these advances in applying knowledge distillation to software analytics, a significant research gap remains in understanding how deeply the student model mimics the teacher’s predictive behavior and internal representations during distillation. Existing approaches primarily rely on standard performance metrics (e.g., accuracy, precision, recall, F1-score, ROUGE, and BLEU-4) to evaluate the performance between student and teacher models. Although these metrics provide insight into task-level performance, they fail to capture how well the student model mimics the teacher’s internal behavior or decision-making patterns. To date, no study has systematically examined whether a student model deeply mimics its teacher’s behavioral characteristics. This limitation underscores a critical gap in current evaluation practices—one that this study addresses by introducing a structured metamorphic testing framework to assess behavioral fidelity between teacher and student models in knowledge distillation.
8 Conclusion
This study investigated the extent to which smaller student models mimic the predictive behavior and internal representations of their larger teacher counterparts in knowledge distillation. Our findings showed that, although student models achieved performance comparable to that of their teachers, their performance deteriorated significantly under adversarial attacks. This degradation underscores their limited ability to deeply mimic the teacher models’ decision-making behavior and internal representations, a limitation that traditional accuracy-based evaluations fail to capture.
To address this limitation, we introduce MetaCompress, a metamorphic testing framework that extends behavioral fidelity assessment beyond accuracy by framing it as a classical software testing problem and applying metamorphic testing principles. Extensive experiments across diverse language models of code, software engineering tasks, and knowledge distillation techniques show that MetaCompress effectively uncovers discrepancies in behavioral fidelity between teacher and student models that are often missed by traditional accuracy-based evaluations. As future work, we plan to integrate MetaCompress into the knowledge distillation pipeline to enable feedback-driven optimization of the distillation process, aiming to reduce the knowledge and representational gaps in student models and ensure behavioral fidelity with their teacher counterparts.
Data Availability
Our code and the corresponding dataset are publicly available to enhance further research [^11] https://doi.org/10.5281/zenodo.16127320.
Acknowledgments
This research is supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grants program, the Canada Foundation for Innovation’s John R. Evans Leaders Fund (CFI-JELF), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).
Declaration of Generative AI and AI-assisted tools
During the preparation of this work, the author(s) used Grammarly Grammarly and ChatGPT OpenAI2025 to find grammatical mistakes and improve sentence clarity/ presentation. After using these tools/services, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v2

@@
 1 Introduction
-Transformer-based vaswani2017attention language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, including clone detection arshad2022codebert, code summarization ahmed2024automatic, vulnerability prediction ding2024vulnerability, and code search chen2024code.
-However, despite their impressive capabilities, the adoption of these models remains limited, particularly in consumer-grade devices (e.g., laptops) and latency-sensitive applications, due to their substantial computational demands, slow inference speed, and considerable energy consumption and carbon footprint shi2022compressing, schwartz2020green.
+Transformer-based Vaswani et al.
+[2017] language models of code have achieved state-of-the-art performance across a wide range of software analytics tasks, including clone detection Arshad et al.
+[2022], code summarization Ahmed et al.
+[2024], vulnerability prediction Ding et al.
+[2024], and code search Chen et al.
+[2024].
+However, despite their impressive capabilities, the adoption of these models remains limited, particularly in consumer-grade devices (e.g., laptops) and latency-sensitive applications, due to their substantial computational demands, slow inference speed, and considerable energy consumption and carbon footprint Shi et al.
+[2023], Schwartz et al.
+[2020].
 For instance, CodeBERT, a widely used code language model with 125 million parameters and a size of 476 MB, can incur up to 1.5 seconds of latency, consume 0.32 kWh of energy, and emit 0.14 kg of CO2 on consumer-grade laptops.
-[^1] When integrated into IDEs, CodeBERT may be invoked thousands of times daily by a developer, reflecting typical usage patterns hellendoorn2019code.
-To address these practical challenges, the software engineering research community has recently begun exploring knowledge distillation methods hinton2015distilling to develop extremely compact models that reduce model size, lower inference latency, and mitigate environmental impact while maintaining task performance shi2022compressing, shi2024greening, shi2024efficient, wei2023towards, saad2024alpine, d2024compression, chen2025smaller.
+[^1] When integrated into IDEs, CodeBERT may be invoked thousands of times daily by a developer, reflecting typical usage patterns Hellendoorn et al.
+[2019].
+To address these practical challenges, the software engineering research community has recently begun exploring knowledge distillation methods Hinton et al.
+[2015] to develop extremely compact models that reduce model size, lower inference latency, and mitigate environmental impact while maintaining task performance Shi et al.
+[2023, 2024a, 2024b], Wei et al.
+[2023], Saad et al.
+[2024], d’Aloisio et al.
+[2024], Chen et al.
+[2025].
 Knowledge distillation (KD) transfers knowledge from a large teacher model to a smaller student [^2] In this study, the term student model refers to the compressed model obtained through knowledge distillation, while the term teacher model denotes the original uncompressed model.
@@
 In this context, Shi et al.
-shi2024greening introduced Compressor, a genetic algorithm-based approach that compresses code-specialized language models (e.g., CodeBERT feng2020codebert) from 476 MB and 125 million parameters to only 3 MB with minimal loss in accuracy.
+Shi et al.
+[2024a] introduced Compressor, a genetic algorithm-based approach that compresses code-specialized language models (e.g., CodeBERT Feng et al.
+[2020]) from 476 MB and 125 million parameters to only 3 MB with minimal loss in accuracy.
 Although KD effectively reduces computational costs and environmental impact, student model performance is typically evaluated using traditional accuracy-based metrics that assess whether the student model’s task performance on ground-truth labels remains comparable to that of the teacher model.
 However, this form of evaluation offers only a surface-level view and may overlook critical discrepancies in behavioral fidelitythe extent to which a student model deeply mimics the predictive behavior and internal representations of its teacher counterpart.
-We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks yang2022natural, zhang2020generating, zeng2022extensive that rename identifiers based on the contextual semantics of each code snippet.
+We investigate the student model’s ability to mimic the teacher by performing simple yet effective adversarial attacks Yang et al.
+[2022], Zhang et al.
+[2020a], Zeng et al.
+[2022] that rename identifiers based on the contextual semantics of each code snippet.
 The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model.
@@
 It also underscores the need for a nuanced evaluation framework that goes beyond accuracy to assess how deeply a student model mimics the predictive behavior and internal representations of its teacher, providing a more comprehensive understanding of model alignment under knowledge distillation.
-To fill this gap, we introduce MetaCompress, a testing framework inspired by metamorphic testing principles in software engineering chen2018metamorphic, to evaluate the behavioral fidelity of the student model in knowledge distillation.
-In contrast to traditional accuracy-based evaluation, which compares model performance against ground-truth labels, MetaCompress performs a head-to-head comparison of student and teacher model outputs across a suite of behavior-preserving metamorphic relations, enabling a more nuanced assessment of behavioral fidelity.
+To fill this gap, we introduce MetaCompress, a testing framework inspired by metamorphic testing principles in software engineering Chen et al.
+[2018], to evaluate the behavioral fidelity of the student model in knowledge distillation.
+In contrast to traditional accuracy-based evaluation, which compares model performance against ground-truth labels, MetaCompress performs a head-to-head comparison of student and teacher model outputs across a set of behavior-preserving metamorphic relations, enabling a more nuanced assessment of behavioral fidelity.
 We design four metamorphic relations, grounded in strong empirical evidence from manual investigations, that capture discrepancies in the behavioral fidelity of student models, ranging from prediction-level disagreement to divergences in the probability distribution.
-We empirically evaluate MetaCompress on two widely studied software engineering tasks, clone detection and vulnerability prediction, using two popular language models of code, CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert.
-To distill their smaller student models, we apply three state-of-the-art knowledge distillation techniques designed for language models of code: Compressor shi2022compressing, AVATAR shi2024greening, and MORPH panichella2025metamorphic.
+We empirically evaluate MetaCompress on two widely studied software engineering tasks, clone detection and vulnerability prediction, using two popular language models of code, CodeBERT Feng et al.
+[2020] and GraphCodeBERT Guo et al.
+[2020].
+To distill their smaller student models, we apply three state-of-the-art knowledge distillation techniques designed for language models of code: Compressor Shi et al.
+[2023], AVATAR Shi et al.
+[2024a], and MORPH Panichella [2025].
 Our extensive experiments show that MetaCompress can reveal behavioral fidelity discrepancies of up to 62% across four metamorphic relations that traditional accuracy-based metrics fail to detect.
@@
 2.1 Language Models of Code
-Transformer-based vaswani2017attention language models of code have demonstrated remarkable performance across a wide range of software analytics tasks, such as vulnerability prediction, clone detection, code summarization, code completion, code comment generation, and code search feng2020codebert, guo2020graphcodebert, lu2021codexglue, ahmad2021unified, wang2021codet5, devlin2018bert, liu2019roberta, sanh2019distilbert, husain2019codesearchnet, hou2024large.
+Transformer-based Vaswani et al.
+[2017] language models of code have demonstrated remarkable performance across a wide range of software analytics tasks, such as vulnerability prediction, clone detection, code summarization, code completion, code comment generation, and code search Feng et al.
+[2020], Guo et al.
+[2020], Lu et al.
+[2021], Ahmad et al.
+[2021], Wang et al.
+[2021], Devlin [2018], Liu [2019], Sanh [2019], Husain et al.
+[2019], Hou et al.
+[2024].
 These models learn general language representations from large amounts of unlabeled data through a two-step process: (i) pre-training and (ii) fine-tuning.
 During pre-training, the model employs a self-supervised objective to acquire a general understanding of source code structures and patterns.
-Once the training is complete, the model can be fine-tuned by updating a few parameters for specific tasks with smaller datasets, making it more efficient in terms of time and computational resources compared to training a model from scratch devlin2018bert.
-Language models for code can be broadly classified into three categories: (1) encoder-only, (2) decoder-only, and (3) encoder-decoder models zeng2022extensive.
-Encoder-only models (e.g., CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert) process the entire input simultaneously and are primarily used for discriminative tasks, such as classification (e.g., vulnerability prediction and clone detection).
-Decoder-only models (e.g., CodeGPT lu2021codexglue, StarCoder lozhkov2024starcoder, and Codex chen2021evaluating) generate output token by token and are typically employed in generative tasks, such as code generation, code completion, and summarization.
-Encoder-decoder models (e.g., PLBART ahmad2021unified and CodeT5 wang2021codet5) encode the input into a latent representation and subsequently decode it into an output sequence, making them suitable for translation-style tasks such as code summarization and programming language translation.
-Among these categories, encoder only models, particularly CodeBERT and GraphCodeBERT, have been widely and effectively applied to various downstream software analytics tasks and model compression studies shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer, as evidenced by prior empirical evaluations lu2021codexglue.
+Once the training is complete, the model can be fine-tuned by updating a few parameters for specific tasks with smaller datasets, making it more efficient in terms of time and computational resources compared to training a model from scratch Devlin [2018].
+Language models for code can be broadly classified into three categories: (1) encoder-only, (2) decoder-only, and (3) encoder-decoder models Zeng et al.
+[2022].
+Encoder-only models (e.g., CodeBERT Feng et al.
+[2020] and GraphCodeBERT Guo et al.
+[2020]) process the entire input simultaneously and are primarily used for discriminative tasks, such as classification (e.g., vulnerability prediction and clone detection).
+Decoder-only models (e.g., CodeGPT Lu et al.
+[2021], StarCoder Lozhkov et al.
+[2024], and Codex Chen et al.
+[2021]) generate output token by token and are typically employed in generative tasks, such as code generation, code completion, and summarization.
+Encoder-decoder models (e.g., PLBART Ahmad et al.
+[2021] and CodeT5 Wang et al.
+[2021]) encode the input into a latent representation and subsequently decode it into an output sequence, making them suitable for translation-style tasks such as code summarization and programming language translation.
+Among these categories, encoder-only models, particularly CodeBERT and GraphCodeBERT, have been widely and effectively applied to various downstream software analytics tasks and model compression studies Shi et al.
+[2023, 2024a, 2024b], Panichella [2025], Liu et al.
+[2026], as evidenced by prior empirical evaluations Lu et al.
+[2021].
 Consequently, these models serve as the primary focus of our study.
 2.2 Model Compression
-A significant barrier to the real-world deployment of language models of code is their substantial computational overhead, high memory demands, and environmental impact (e.g., carbon emissions) schwartz2020green, xu2023survey.
+A significant barrier to the real-world deployment of language models of code is their substantial computational overhead, high memory demands, and environmental impact (e.g., carbon emissions) Schwartz et al.
+[2020], Xu and McAuley [2023].
 To mitigate these challenges and advance sustainability, a widely adopted strategy known as model compression offers a solution that reduces the resource consumption of such models without significantly compromising performance.
@@
 In addition to deployment benefits, model compression improves inference speed and enhances overall resource efficiency.
-Algorithmically, the fundamental model compression techniques include: pruning sanh2020movement, quantization zafrir2019q8bert, knowledge distillation hinton2015distilling, neural architecture search elsken2019neural, weight sharing & hashing chen2015compressing, and low-rank factorization sainath2013low-each method uniquely contributes to model optimization.
+Algorithmically, the fundamental model compression techniques include: pruning Sanh et al.
+[2020], quantization Zafrir et al.
+[2019], knowledge distillation Hinton et al.
+[2015], neural architecture search Elsken et al.
+[2019], weight sharing & hashing Chen et al.
+[2015], and low-rank factorization Sainath et al.
+[2013]-each method uniquely contributes to model optimization.
 Among these, pruning, quantization, and knowledge distillation (KD) have emerged as the most prominent and widely adopted approaches for compressing language models of code in recent software engineering research.
-shi2022compressing, shi2024greening, wei2023towards, d2024compression, chen2025smaller, liu2026pioneer, afrin2025quantization.
+Shi et al.
+[2023, 2024a], Wei et al.
+[2023], d’Aloisio et al.
+[2024], Chen et al.
+[2025], Liu et al.
+[2026], Afrin et al.
+[2025].
 While pruning and quantization primarily remove redundant parameters or reduce numerical precision to achieve compactness, their ability to produce highly compact models is inherently limited, as aggressive compression often results in substantial degradation of representational capacity.
@@
 2.3 Metamorphic Testing
-Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem chen2018metamorphic, zhang2014search.
+Metamorphic testing (MT) is a property-based testing paradigm widely adopted for validating software systems in the absence of a test oracle or when one is difficult to define, a challenge commonly referred to as the oracle problem Chen et al.
+[2018], Zhang et al.
+[2014].
 Rather than verifying the correctness of individual outputs in isolation, MT addresses this issue by evaluating whether the program under test exhibits expected relationships between outputs across a set of well-defined metamorphic relations (MRs).
@@
 These relations are derived from the problem domain’s inherent semantics and define how the output should change (or remain consistent) when the input is systematically modified.
-This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed yang2025hallucination, xie2011testing, chen2020metamorphic, xiao2022metamorphic, segura2018metamorphic.
+This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed Yang et al.
+[2025], Xie et al.
+[2011], Chen et al.
+[2020], Xiao et al.
+[2022], Segura et al.
+[2018].
 In this study, we adapt MT to evaluate how deeply a student model produced via KD mimics its teacher model’s behavior.
@@
 The concept of adversarial attack was first introduced by Zegedy et al.
-szegedy2013intriguing, who demonstrated that state-of-the-art image classifiers can be misled by carefully designed pixel-level modifications that remain imperceptible to human observers.
+Szegedy et al.
+[2013], who demonstrated that state-of-the-art image classifiers can be misled by carefully designed pixel-level modifications that remain imperceptible to human observers.
 Formally, given a classifier $f:X\rightarrow Y$ that correctly predicts the label $y_{truth}\in Y$ for an input $x\in X$, an adversarial attack generates a perturbed version $x^{\prime}$ such that $f(x^{\prime})\neq f(x)=y_{truth}$.
 In continuous domains such as images, producing visually indistinguishable adversarial examples is relatively straightforward.
-However, generating adversarial examples for discrete domains like source code is considerably more complex due to the rigid syntactic and grammatical structure of programming languages zhang2022towards, zhang2020generating.
+However, generating adversarial examples for discrete domains like source code is considerably more complex due to the rigid syntactic and grammatical structure of programming languages Zhang et al.
+[2022, 2020a].
 As a result, adversarial examples designed to attack language models of code must meet three fundamental requirements.
@@
 A model is considered adversarially robust if it continues to generate consistent, accurate outputs despite deliberate input modifications intended to mislead it.
-In the context of machine learning, this concept reflects a model’s resilience to adversarial attacks that exploit its sensitivity to small perturbations in the input space szegedy2013intriguing, zhang2022towards.
+In the context of machine learning, this concept reflects a model’s resilience to adversarial attacks that exploit its sensitivity to small perturbations in the input space Szegedy et al.
+[2013], Zhang et al.
+[2022].
 2.6 Kullback–Leibler Divergence
-The Kullback-Leibler (KL) divergence kullback1951information is a measure of how one probability distribution $Q$ diverges from a reference probability distribution $P$.
+The Kullback-Leibler (KL) divergence Kullback and Leibler [1951] is a measure of how one probability distribution $Q$ diverges from a reference probability distribution $P$.
 For discrete distributions defined over the same support, it is mathematically expressed as:
@@
 3.1.1 Knowledge Distillation Approaches
-In recent years, several knowledge distillation (KD) approaches have been proposed for language models of code, including Compressor shi2022compressing, AVATAR shi2024greening, SODA chen2025smaller, MORPH panichella2025metamorphic, and PIONEER liu2026pioneer.
+In recent years, several knowledge distillation (KD) approaches have been proposed for language models of code, including Compressor Shi et al.
+[2023], AVATAR Shi et al.
+[2024a], SODA Chen et al.
+[2025], MORPH Panichella [2025], and PIONEER Liu et al.
+[2026].
 In this study, we select Compressor, AVATAR, and MORPH as our KD approaches to generate student models, as the replication package for SODA is not fully available [^4] https://shorturl.at/Se0f7.
@@
 $\mathcal{L}=-\frac{1}{n}\sum_{i}^{n}softmax(\frac{p_{i}}{T})\text{log}\Bigl(softmax(\frac{q_{i}}{T})\Bigl)T^{2}$ (2)
-In “(2)”, $p_{i}$ and $q_{i}$ represent the teacher and student model logits, respectively, with $T$ denoting the temperature parameter controlling the softmax distribution hinton2015distilling.
+In “(2)”, $p_{i}$ and $q_{i}$ represent the teacher and student model logits, respectively, with $T$ denoting the temperature parameter controlling the softmax distribution Hinton et al.
+[2015].
 While the teacher logits remain fixed, the student logits are updated throughout the training process.
@@
 To minimize experimental bias in evaluating whether the student model deeply mimics the teacher’s predictive behavior and internal representations during knowledge distillation, we adopt the same models and datasets used in the knowledge distillation–based approaches described in Section 3.1.1.
-Specifically, we select CodeBERT feng2020codebert and GraphCodeBERT guo2020graphcodebert models and two well-studied downstream tasks: clone detection and vulnerability prediction, which serve as standard benchmarks in knowledge distillation research for software engineering shi2022compressing, shi2024greening, shi2024efficient, panichella2025metamorphic, liu2026pioneer.
+Specifically, we select CodeBERT Feng et al.
+[2020] and GraphCodeBERT Guo et al.
+[2020] models and two well-studied downstream tasks: clone detection and vulnerability prediction, which serve as standard benchmarks in knowledge distillation research for software engineering Shi et al.
+[2023, 2024a, 2024b], Panichella [2025], Liu et al.
+[2026].
 Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent.
-To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset svajlenko2014towards provided by Wang et al.
-wang2020detecting.
+To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset Svajlenko et al.
+[2014] provided by Wang et al.
+Wang et al.
+[2020].
 The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing.
@@
 We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al.
-zhou2019devign.
+Zhou et al.
+[2019].
 The dataset comprises 27,318 functions sourced from two widely used open-source C libraries, FFmpeg and Qemu.
 Each function is manually annotated to indicate whether it contains vulnerabilities.
-We use the CodeXGLUE benchmark splits lu2021codexglue, including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
-CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa liu2019roberta.
+We use the CodeXGLUE benchmark splits Lu et al.
+[2021], including training, development, and test sets with 27,318, 2,732, and 2,732 examples, respectively.
+CodeBERT is a pre-trained language model that shares the same architecture and a similar pre-training strategy as RoBERTa Liu [2019].
 It consists of twelve Transformer layers, each with 12-headed multi-head attention mechanisms and a hidden size of 768.
 CodeBERT is pre-trained using two objectives: masked language modeling, which predicts masked tokens, and replaced token detection, which identifies substituted tokens in the input.
-The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset husain2019codesearchnet.
+The model is trained on bimodal data comprising natural language–code pairs across six programming languages from the CodeSearchNet dataset Husain et al.
+[2019].
 GraphCodeBERT shares CodeBERT’s architecture but incorporates data-flow information from source code during pre-training.
@@
 Following the recommendation of Du et al.
-du2023extensive, we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code yang2022natural, zhang2020generating.
+Du et al.
+[2023], we adopt black-box attacks, as they rely solely on model outputs and are widely used for evaluating language models of code Yang et al.
+[2022], Zhang et al.
+[2020a].
 The perturbations introduced through these attacks are deliberately subtle to maintain the syntactic and semantic integrity of the source code, as discussed in Section 2.4, while inducing mispredictions.
-To this end, we employ three established black-box techniques: ALERT yang2022natural, Metropolis Hastings Modifier (MHM) zhang2020generating, and WIR-Random zeng2022extensive, each of which performs identifier substitution to assess model robustness without altering code functionality.
-ALERT yang2022natural considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier.
-MHM zhang2020generating utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration.
-WIR-Random zeng2022extensive ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates.
-It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications yang2022natural, zhang2020generating.
+To this end, we employ three established black-box techniques: ALERT Yang et al.
+[2022], Metropolis Hastings Modifier (MHM) Zhang et al.
+[2020a], and WIR-Random Zeng et al.
+[2022], each of which performs identifier substitution to assess model robustness without altering code functionality.
+ALERT Yang et al.
+[2022] considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier.
+MHM Zhang et al.
+[2020a] utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration.
+WIR-Random Zeng et al.
+[2022] ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates.
+It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications Yang et al.
+[2022], Zhang et al.
+[2020a].
 3.1.4 Experimental Settings
-We fine-tune CodeBERT using the hyperparameter settings and training pipeline provided by the CodeXGLUE benchmark lu2021codexglue, while for GraphCodeBERT, we adopt the settings specified in its original paper guo2020graphcodebert.
+We fine-tune CodeBERT using the hyperparameter settings and training pipeline provided by the CodeXGLUE benchmark Lu et al.
+[2021], while for GraphCodeBERT, we adopt the settings specified in its original paper Guo et al.
+[2020].
 To ensure consistency across our experiments, we adopt the data-splitting strategy used in knowledge distillation–based approaches such as MORPH and AVATAR, which divide the training data into two equal halves: one for fine-tuning and the other for knowledge distillation.
 We apply the same fine-tuning split to all models.
-Despite being trained on only half of the original dataset, all models achieve performance comparable to the state-of-the-art results reported in the CodeXGLUE benchmark lu2021codexglue, the GraphCodeBERT paper guo2020graphcodebert, and knowledge distillation–based compression studies such as Compressor, AVATAR, and MORPH.
+Despite being trained on only half of the original dataset, all models achieve performance comparable to the state-of-the-art results reported in the CodeXGLUE benchmark Lu et al.
+[2021], the GraphCodeBERT paper Guo et al.
+[2020], and knowledge distillation–based compression studies such as Compressor, AVATAR, and MORPH.
 3.1.5 Evaluation Metrics
-Following the evaluation protocols of CodeXGLUE lu2021codexglue, Yang et al.
-yang2022natural, and Shi et al.
-shi2022compressing, shi2024greening, we use accuracy as the primary metric to evaluate model performance on downstream tasks.
+Following the evaluation protocols of CodeXGLUE Lu et al.
+[2021], Yang et al.
+Yang et al.
+[2022], and Shi et al.
+Shi et al.
+[2023, 2024a], we use accuracy as the primary metric to evaluate model performance on downstream tasks.
 To evaluate the performance of both teacher and student models under adversarial conditions, we measure the effectiveness of the applied attack methods.
 A higher attack effectiveness indicates greater model vulnerability to adversarial perturbations.
-We employ the Attack Success Rate (ASR) metric yang2022natural to measure the effectiveness of an attack method.
+We employ the Attack Success Rate (ASR) metric Yang et al.
+[2022] to measure the effectiveness of an attack method.
 ASR measures the proportion of adversarial samples for which the model fails to produce a correct prediction after the attack and is formally defined as follows:
@@
 Because the inputs for the selected software analytics tasks are source code, we employ four established metrics from Du et al.
-du2023extensive: Identifier Change Rate (ICR), Token Change Rate (TCR), Average Edit Distance (AED), and Average Code Similarity (ACS), to measure the quality of the adversarial examples.
+Du et al.
+[2023]: Identifier Change Rate (ICR), Token Change Rate (TCR), Average Edit Distance (AED), and Average Code Similarity (ACS), to measure the quality of the adversarial examples.
 For a set of $m$ adversarial examples, each $i$-th code snippet contains $k_{i}$ identifiers, among which $n_{i}$ are modified to generate the adversarial example. The ICR is computed as follows:
@@
 Although a slight degradation is observed in certain cases—for instance, the MHM attack shows comparatively higher AED values—the overall differences remain limited.
-We further perform the Friedman test friedman1937use to determine whether the three attack techniques (ALERT, MHM, and WIR-Random) generate adversarial examples of significantly different quality for each model variant of CodeBERT and GraphCodeBERT.
+We further perform the Friedman test Friedman [1937] to determine whether the three attack techniques (ALERT, MHM, and WIR-Random) generate adversarial examples of significantly different quality for each model variant of CodeBERT and GraphCodeBERT.
 A $p$-value below 0.05 indicates that the differences in the quality of adversarial examples produced by the three attack techniques are statistically significant.
@@
 To investigate this, we compare the quality of adversarial examples generated by each attack technique for the teacher and student models across all the evaluation metrics.
-The Wilcoxon signed-rank test woolson2007wilcoxon is used to assess the statistical significance of these differences.
+The Wilcoxon signed-rank test Woolson [2007] is used to assess the statistical significance of these differences.
 The comparison between the Teacher and AVATAR models ($p=0.13$) and between the Teacher and MORPH models ($p=0.85$) indicates that the quality of adversarial examples remains statistically comparable, with no significant differences observed in either case.
@@
 Motivated by this observation, we introduce MetaCompress, a metamorphic testing framework for evaluating the behavioral fidelity of compressed models by framing the task as a classical software testing problem and applying a metamorphic testing strategy.
-The framework identifies key behavioral invariants that should be preserved after model compression and formalizes a set of output-based metamorphic relations that serve as oracle-free tests to assess behavioral fidelity between teacher and student models.
+The framework identifies key behavioral invariants that should be preserved after knowledge distillation and formalizes a set of output-based metamorphic relations that serve as oracle-free tests to assess behavioral fidelity between teacher and student models.
 Figure 3 presents the overall design and workflow of the proposed MetaCompress framework.
@@
 That is, whether the student model produces outputs that are behaviorally consistent with those of the teacher model across all test cases.
-In traditional software engineering, techniques such as regression testing wong1997study, mutation testing jia2010analysis, and metamorphic testing chen2018metamorphic are commonly employed to assess the reliability of modified systems.
+In traditional software engineering, techniques such as regression testing Wong et al.
+[1997], mutation testing Jia and Harman [2010], and metamorphic testing Chen et al.
+[2018] are commonly employed to assess the reliability of modified systems.
 However, evaluating language models of code presents unique challenges compared to traditional software systems.
@@
 These relations act as relational oracles, enabling us to verify whether the teacher and student models produce consistent outputs to the same inputs.
-Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models xie2020mettle, sun2024fairness, kwan2024mt, zhang2020machine, wu2025detecting, yang2025hallucination, demonstrating its versatility and effectiveness in oracle-free evaluation scenarios.
+Finally, metamorphic testing has been widely adopted for evaluating machine learning–based software systems and large language models Xie et al.
+[2020], Sun et al.
+[2024], Kwan et al.
+[2024], Zhang et al.
+[2020b], Wu et al.
+[2025], Yang et al.
+[2025], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios.
 Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
@@
 To quantify the difference between the two probability distributions, we use the Kullback–Leibler divergence, as defined in “(1)”.
-Given a tolerance $\delta=0.5$ (Inspired by sigmoid-based methods for binary classification lecun2015deep), we then count the test instances for which MR2 holds:
+Given a tolerance $\delta=0.5$ (Inspired by sigmoid-based methods for binary classification LeCun et al.
+[2015]), we then count the test instances for which MR2 holds:
 $\text{Loyalty}_{\text{prob}}=\frac{1}{|X|}\sum_{x\in X}\mathbf{1}\!\left[D_{\mathrm{KL}}\!\bigl(P_{T}(\cdot\mid x),P_{S}(\cdot\mid x)\bigr)\leq\delta\right]\qquad$ (9)
@@
 Poor calibration, where predicted probabilities do not reflect true likelihoods, can lead to over- or under-confidence, which adversely affects downstream decision-making.
-Prior studies shang2024enhancing, ji2024beware on vision and language models demonstrate that compression techniques such as quantization can significantly distort model calibration, even when accuracy remains unaffected.
+Prior studies Shang et al.
+[2024], Ji et al.
+[2024] on vision and language models demonstrate that compression techniques such as quantization can significantly distort model calibration, even when accuracy remains unaffected.
 These findings motivate the design of this MR to assess behavioral fidelity through the lens of calibration alignment.
@@
 Ba et al.
-ba2014deep stated that ”If the student model learns to mimic the teacher model perfectly, it makes exactly the same predictions and mistakes as the teacher model.” However, achieving such mimicry is extremely challenging when there exists a substantial gap in network capacity wang2021knowledge.
+Ba and Caruana [2014] stated that ”If the student model learns to mimic the teacher model perfectly, it makes exactly the same predictions and mistakes as the teacher model.” However, achieving such mimicry is extremely challenging when there exists a substantial gap in network capacity Wang and Yoon [2021].
 For instance, the knowledge distillation technique Compressor reduces the size of the CodeBERT model from 476 MB to only 3 MB.
@@
     Wang et al.
-wang2025empirical empirically demonstrated that the same student model distilled from a teacher performs significantly better than when trained using the standard fine-tuning approach on the same task.
+Wang et al.
+[2025] empirically demonstrated that the same student model distilled from a teacher performs significantly better than when trained using the standard fine-tuning approach on the same task.
 Since the teacher model typically possesses much greater network capacity, it can learn more complex decision boundaries and have greater robustness against adversarial attacks.
@@
 A potential threat to internal validity may arise if differences in performance between the student and teacher models are influenced by inconsistent training procedures, fine-tuning configurations, or data handling, rather than the knowledge distillation approaches themselves.
-To mitigate this threat, we applied uniform training protocols and hyperparameter settings across all models, both student and teacher, following the configurations reported in prior knowledge distillation studies shi2022compressing, shi2024greening, panichella2025metamorphic.
+To mitigate this threat, we applied uniform training protocols and hyperparameter settings across all models, both student and teacher, following the configurations reported in prior knowledge distillation studies Shi et al.
+[2023, 2024a], Panichella [2025].
 Additionally, we used identical data splits for fine-tuning and evaluation to ensure consistency and fairness in comparison.
@@
 7 Related Work
-A substantial body of research on model compression has emerged in the fields of computer vision and natural language processing (NLP), aiming to develop large language models that are energy efficient, environmentally sustainable, and optimized for memory usage xu2023survey, zhu2024survey, xu2025resource, xu2021survey, gordon2020compressing, xu2021beyond, ye2019adversarial, du2021robustness, gourtani2024improving, zhu2022safety, sanh2019distilbert, sun2019patient, jiao2019tinybert, buciluǎ2006model, jiang2023lion, zhang2018structadmm, tang2019distilling, xu2020bert.
-To this end, several techniques, including pruning sanh2020movement, quantization zafrir2019q8bert, and knowledge distillation (KD) hinton2015distilling, have been widely explored.
+A substantial body of research on model compression has emerged in the fields of computer vision and natural language processing (NLP), aiming to develop large language models that are energy efficient, environmentally sustainable, and optimized for memory usage Xu and McAuley [2023], Zhu et al.
+[2024], Xu et al.
+[2025, 2021b], Gordon et al.
+[2020], Xu et al.
+[2021a], Ye et al.
+[2019], Du et al.
+[2021], Gourtani and Meratnia [2024], Zhu et al.
+[2022], Sanh [2019], Sun et al.
+[2019], Jiao et al.
+[2019], buciluǎ2006model, Jiang et al.
+[2023], Zhang et al.
+[2018], Tang et al.
+[2019], Xu et al.
+[2020].
+To this end, several techniques, including pruning Sanh et al.
+[2020], quantization Zafrir et al.
+[2019], and knowledge distillation (KD) Hinton et al.
+[2015], have been widely explored.
 More recently, these compression techniques have been extended to software analytics to improve the efficiency of language models of code.
 Wei et al.
-wei2023towards evaluated quantized models on code generation tasks, examining accuracy, resource consumption, and environmental impact, and found that quantization can substantially improve model efficiency with minimal degradation in accuracy or robustness when applied under suitable conditions.
+Wei et al.
+[2023] evaluated quantized models on code generation tasks, examining accuracy, resource consumption, and environmental impact, and found that quantization can substantially improve model efficiency with minimal degradation in accuracy or robustness when applied under suitable conditions.
 Saad et al.
-saad2024alpine introduced ALPINE, a pruning method that is independent of programming language and effectively reduces the computational cost of code-language models.
+Saad et al.
+[2024] introduced ALPINE, a pruning method that is independent of programming language and effectively reduces the computational cost of code-language models.
 Aloisio et al.
-d2024compression conducted an empirical study on how different compression strategies, such as pruning, quantization, and knowledge distillation, affect the performance and resource efficiency of the CodeBERT model across three software analytics tasks: code summarization, vulnerability detection, and code search.
+d’Aloisio et al.
+[2024] conducted an empirical study on how different compression strategies, such as pruning, quantization, and knowledge distillation, affect the performance and resource efficiency of the CodeBERT model across three software analytics tasks: code summarization, vulnerability detection, and code search.
 Afrin et al.
-afrin2025quantization empirically examined the effects of quantizing large code models beyond functional correctness, focusing on qualitative aspects and static properties of the generated code.
+Afrin et al.
+[2025] empirically examined the effects of quantizing large code models beyond functional correctness, focusing on qualitative aspects and static properties of the generated code.
 While pruning and quantization improve efficiency, their compression capacity is limited, as excessive compression often harms model performance.
@@
 Shi et al.
-shi2022compressing proposed Compressor, a KD-based method evaluated on CodeBERT and GraphCodeBERT for two code understanding tasks: vulnerability prediction and clone detection.
+Shi et al.
+[2023] proposed Compressor, a KD-based method evaluated on CodeBERT and GraphCodeBERT for two code understanding tasks: vulnerability prediction and clone detection.
 Experimental results show that Compressor significantly reduces inference time and model size, down to 3 MB, while preserving model effectiveness.
 In a follow-up study, Shi et al.
-shi2024greening introduced AVATAR, another KD-based approach designed to mitigate energy consumption and carbon emissions, while also maintaining model effectiveness and reducing inference latency.
-Building on AVATAR, MORPH panichella2025metamorphic was proposed as a KD method that combines metamorphic testing with many-objective optimization to enhance model robustness on metamorphic code, as well as accuracy, size, and efficiency.
+Shi et al.
+[2024a] introduced AVATAR, another KD-based approach designed to mitigate energy consumption and carbon emissions, while also maintaining model effectiveness and reducing inference latency.
+Building on AVATAR, MORPH Panichella [2025] was proposed as a KD method that combines metamorphic testing with many-objective optimization to enhance model robustness on metamorphic code, as well as accuracy, size, and efficiency.
 Experiments on CodeBERT and GraphCodeBERT showed that MORPH achieved up to 73% greater robustness and 38% greater efficiency than AVATAR, without sacrificing accuracy or increasing model size.
@@
 Recently, Liu et al.
-liu2026pioneer proposed a KD-based approach, PIONEER, that enhances the robustness of compressed code models while preserving accuracy.
+Liu et al.
+[2026] proposed a KD-based approach, PIONEER, that enhances the robustness of compressed code models while preserving accuracy.
 Using a solver-constrained genetic algorithm and robustness evaluation with ALERT-generated examples, it achieved an 87.54% robustness gain with only a 1.67% drop in accuracy across CodeBERT, GraphCodeBERT, and CodeT5.
 Recently, Wang et al.
-wang2025empirical empirically evaluated the effectiveness of logit-based and feature-based KD approaches for code understanding tasks.
+Wang et al.
+[2025] empirically evaluated the effectiveness of logit-based and feature-based KD approaches for code understanding tasks.
 Their study showed that KD consistently outperforms fine-tuning on code understanding tasks, with feature-based methods and code-specific teachers yielding the best results.
@@
 Evans Leaders Fund (CFI-JELF), and by the industry-stream NSERC CREATE in Software Analytics Research (SOAR).
-Declaration of Generative AI and AI-assisted tools
-During the preparation of this work, the author(s) used Grammarly Grammarly and ChatGPT OpenAI2025 to find grammatical mistakes and improve sentence clarity/ presentation.
+Declaration of generative AI and AI-assisted technologies in the manuscript preparation process
+During the preparation of this work, the author(s) used Grammarly Inc.
+[2025] and ChatGPT OpenAI [2025] to find grammatical mistakes and improve sentence clarity/ presentation.
 After using these tools/services, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v3

@@
 [2020a], Zeng et al.
-[2022] that rename identifiers based on the contextual semantics of each code snippet.
+[2022], Tian et al.
+[2023].
+These attacks involve renaming identifiers and applying semantic-preserving code transformations informed by the contextual semantics of each code snippet.
 The empirical results reveal that although the student models achieve performance comparable to the teacher before adversarial perturbations, their performance degrades substantially under adversarial attacks, with up to 285% greater performance loss than the teacher model.
@@
 This paradigm is beneficial for testing complex systems such as machine learning models, scientific computing software, or simulation tools, where correctness is difficult to assert for individual outputs but consistency across related cases can be meaningfully assessed Yang et al.
-[2025], Xie et al.
+[2025a], Xie et al.
 [2011], Chen et al.
@@
 [2026].
-Clone detection: (CD) This task aims to determine whether two given code snippets are syntactically or semantically equivalent.
+Clone detection (CD): This task aims to determine whether two given code snippets are syntactically or semantically equivalent.
 To streamline the fine-tuning process, we adopt the filtered BigCloneBench dataset Svajlenko et al.
@@
 The final dataset includes 90,102 training examples, along with 4,000 examples each for validation and testing.
-Vulnerability prediction: (VP) This task aims to determine whether a given code snippet contains a vulnerability.
+Vulnerability prediction (VP): This task aims to determine whether a given code snippet contains a vulnerability.
 We conduct our experiments using the Devign [^5] https://sites.google.com/view/devign dataset, introduced by Zhou et al.
@@
 [2022], each of which performs identifier substitution to assess model robustness without altering code functionality.
-ALERT Yang et al.
-[2022] considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier.
-MHM Zhang et al.
-[2020a] utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration.
-WIR-Random Zeng et al.
-[2022] ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates.
+To further enhance robustness evaluation, we also employ CODA Tian et al.
+[2023], a black-box attack technique that combines identifier renaming with semantic-preserving code transformations.
+This enables assessment of model robustness under a wider range of functionally equivalent program variations.
+ALERT considers both operational and natural semantics, using greedy and genetic algorithms to select substitutions, with 30 candidates per identifier.
+MHM utilizes Metropolis-Hastings sampling for iterative identifier substitution, with a maximum of 100 iterations and 30 candidates per iteration.
+WIR-Random ranks identifiers based on their impact on model predictions and replaces them sequentially with randomly chosen candidates.
+In contrast to these identifier-centric techniques, CODA utilizes code differences between a target input and reference inputs with divergent predictions to guide adversarial example generation.
+It combines identifier renaming with semantic-preserving structural transformations.
+By considering both the identifier level and structural differences, CODA generates functionally equivalent yet more diverse adversarial examples.
 It is important to note that only instances correctly classified before perturbation are included, ensuring that the robustness evaluation is based exclusively on valid predictions and that any observed degradation in model performance is caused solely by adversarial perturbations rather than pre-existing misclassifications Yang et al.
@@
 Figure 2: Attack success rate metric values across adversarial attack techniques, models, tasks, and knowledge distillation methods.
+To further assess the robustness of teacher and student models to a broader class of semantic-preserving adversarial perturbations, we conduct additional experiments using the CODA attack.
+Figure 3 presents the Attack Success Rate values under CODA.
+The results align with the trends observed for ALERT, MHM, and WIR-Random attacks.
+Across both tasks and model architectures, student models trained via knowledge distillation generally achieve higher ASR values than their teacher counterparts, indicating reduced robustness to CODA-induced perturbations.
+This finding applies to both AVATAR and MORPH students, confirming that the previously observed robustness degradation is not restricted to identifier-level attacks.
+Figure 3: Attack success rate metric values across CODA adversarial attack, models, tasks, and knowledge distillation methods.
 Taken together, these findings suggest that student models distilled from their teachers, particularly those produced by AVATAR and, to a lesser extent, by MORPH, are more susceptible to adversarial attacks than their teacher counterparts.
@@
 [2025], Yang et al.
-[2025], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios.
+[2025a], demonstrating its versatility and effectiveness in oracle-free evaluation scenarios.
 Building on this foundation, MetaCompress adopts a metamorphic testing–based approach to evaluate behavioral fidelity of student models, framing it as an oracle-free testing problem.
@@
 5 Discussion
+5.1 Behavioral Fidelity and Knowledge Gap in Knowledge Distillation
 Ba et al.
@@
 These insights open promising avenues for advancing knowledge distillation methods that achieve greater alignment between student and teacher models in terms of predictive behavior, internal representations, generalizability, and robustness.
+5.2 Practical Consideration and Deployment Cost
+From a practical standpoint, evaluating the behavioral fidelity does not impose significant overhead.
+MetaCompress is designed exclusively as an evaluation framework to determine whether a distilled student model deeply mimics its teacher prior to deployment.
+In standard knowledge distillation workflows, both the teacher and student models are evaluated on held-out test data to assess predictive performance, which requires forward passes through both models.
+MetaCompress leverages the probability distributions and related outputs generated during this routine evaluation as byproducts, without requiring additional model executions.
+Therefore, even for teams that frequently conduct model distillation, MetaCompress introduces only negligible computational overhead beyond lightweight post-processing.
+This approach is both practical and scalable for diagnosing behavioral fidelity gaps and ensuring that compressed student models are robust and reliable for real-world deployment.
+5.3 Actionable Insights for Practitioners and Researchers
+Although MetaCompress reports violation rates as diagnostic measures of behavioral fidelity divergence, these values should not be interpreted in isolation.
+Instead, they serve as actionable indicators that inform decision-making throughout the model development and deployment pipeline.
+5.3.1 Insights for Practitioners
+  1. 1.
+    The appropriate response to an observed metamorphic relation (MR) violation rate depends on the intended deployment context and the acceptable level of risk.
+For example, a student model with a low violation rate (e.g., below 10%) may be considered sufficiently stable for use in non-critical applications, such as software clone detection.
+In contrast, for security-critical tasks (e.g., vulnerability detection) or adversarially exposed environments, practitioners should exercise greater caution and avoid deploying highly compact models.
+  2. 2.
+    The acceptable MR violation rate is context-dependent and varies according to the risk tolerance of the specific application.
+It is recommended that practitioners integrate violation-rate checks into continuous integration and deployment pipelines to serve as a gating criterion for model release.
+Models exhibiting low violation rates may be suitable for deployment in non-critical applications.
+In contrast, models designed for security-sensitive or adversarial environments should adhere to more stringent thresholds or be supplemented with additional safeguards, such as adversarial defenses or post-deployment monitoring, prior to release.
+  3. 3.
+    When violation rates indicate substantial behavioral divergence between teacher and student models within a specific application context, practitioners may consider revisiting the selected student architecture or knowledge distillation strategy.
+Adopting alternative student configurations, improved supervision mechanisms, or multi-teacher distillation approaches may mitigate metamorphic relation violations and reduce the knowledge gap, thereby enhancing behavioral fidelity prior to deployment.
+5.3.2 Insights for Researchers
+  1. 1.
+    From a research perspective, MR violation rates provide a quantitative signal for diagnosing knowledge transfer quality beyond conventional accuracy metrics.
+Since student models exhibit lower adversarial robustness than their teacher counterparts, the soft knowledge transferred from the teacher to the student does not adequately reflect the teacher’s decision boundary; hence, the researcher should focus on improving the quality of the transferred knowledge.
+  2. 2.
+    Yang et al.
+Yang et al.
+[2025b] demonstrated that Mixture of Experts (MoE) frameworks enhance software vulnerability detection relative to single-model approaches.
+Building on this finding, subsequent research could investigate integrating MoE architectures into knowledge distillation to extract complementary knowledge from multiple specialized teachers, rather than relying on a single generalist model.
+  3. 3.
+    In this study, the selected knowledge distillation approaches focus on aligning the final-layer outputs of the teacher and student models as the primary training objective.
+While this design choice is widely adopted, future research may explore incorporating intermediate-layer alignment as an additional or complementary objective to further reduce the knowledge gap between teacher and student models and potentially improve behavioral fidelity.
+In summary, the insights above suggest that behavioral fidelity divergence and metamorphic relation violation rates should be interpreted as decision-support metrics rather than pass–fail criteria.
+By making decision-boundary instability observable and measurable, MetaCompress enables developers to make informed follow-up decisions—such as pursuing architectural refinements or accepting a model based on application-specific risk tolerance—and encourages researchers in designing more robust and reliable knowledge distillation techniques.
 6 Threats to Validity